### v1

# MEEG and AT-DGNN: Advancing EEG Emotion Recognition with Music and Graph Learning

Abstract
Recent advances in neuroscience have elucidated the crucial role of coordinated brain region activities during cognitive tasks. To explore this complexity, we introduce the MEEG dataset—a comprehensive multi-modal music-induced electroencephalogram (EEG) dataset and the Attention-based Temporal learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition. The MEEG dataset captures a wide range of emotional responses to music, enabling an in-depth analysis of brainwave patterns in musical contexts. The AT-DGNN combines an attention-based temporal learner with a dynamic graph neural network (DGNN) to accurately model the local and global graph dynamics of EEG data across varying brain network topologies. Our evaluations show that AT-DGNN achieves superior performance, with accuracies (ACC) of 83.06% in arousal and 85.31% in valence, outperforming state-of-the-art (SOTA) methods on the MEEG dataset. Comparative analyses with traditional datasets like DEAP highlight the effectiveness of our approach and underscore the potential of music as a powerful medium for emotion induction. This study not only advances our understanding of the brain’s emotional processing but also enhances the accurcy of emotion recognition technologies in brain-computer interfaces (BCI), leveraging both graph-based learning and the emotional impact of music. The source code and dataset are available at https://github.com/xmh1011/AT-DGNN.
Index Terms—Electroencephalogram (EEG), Emotion recognition, Dynamic graph neural networks (DGNNs).
I Introduction
Emotion, a complex physiological and psychological phenomenon, plays a critical role in interpreting genuine reactions to daily interactions [1]. In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2]. These models conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
BCI technology, which decodes brain signals into commands for external devices, predominantly employs EEG signals [3]. This technology processes EEG signals through stages of preprocessing, feature extraction, and classification, thereby translating thoughts into actionable commands and enhancing human-machine interaction [4]. EEG signals, reflecting cortical voltage fluctuations, are crucial for emotion recognition due to their objectivity, high temporal resolution, affordability, and rapid data acquisition [5].
EEG signals have long been pivotal in neuroscience research [6]. In 2011, Koelstra et al. established the DEAP emotion dataset by collecting physiological signals associated with emotional responses [7]. Subsequently, in 2015, Zheng et al. developed the SEED dataset, providing well-annotated EEG signals from multiple subjects [8]. Both datasets have been instrumental in facilitating the development and evaluation of computational models for emotion analysis. Over recent decades, diverse machine learning and signal processing techniques have been utilized for emotion recognition from physiological signals.
A significant advancement came in 2018 when Lawhern et al. introduced EEGNet, a compact convolutional neural network (CNN) designed for efficient EEG signal processing [9]. Further advancements were made by Schirrmeister et al., who enhanced EEG decoding using deep learning with CNNs and provided sophisticated visualization tools to better comprehend brain activities [10]. Temporal convolutional networks (TCNs) have also become a formidable alternative for real-time BCI applications. In 2020, Ingolfsson et al. developed EEG-TCNet, demonstrating high accuracy in motor imagery tasks and suggesting that TCNs could surpass traditional methods in specific BCI applications [11]. This was supported by Musallam et al., who highlighted the versatility and efficiency of TCNs in complex BCI tasks through the fusion of TCN for motor imagery classification [12].
All the aforementioned research treated EEG signals as two-dimensional time-series data, with the dimensions being channels and time [13]. Channels refer to EEG electrodes, which are strategically placed on the scalp to record brain activity from various functional areas according to the 10-20 system established by the International Federation of Clinical Neurophysiology [3]. In recent years, EEG data has been treated as a graph where electrodes are positioned in a two-dimensional layout that reflects their physical placement on the scalp [14]. In this context, the raw data or extracted features from each electrode introduce a third dimension to this spatial map [15].
In 2023, Ding et al. [16] introduced the Local-Global-Graph network (LGGNet), which processed EEG data in an image-like format and has shown enhanced classification performance, although it was primarily designed for multi-task classification and didn’t significantly improve emotion recognition accuracy compared to other models. In terms of EEG datasets, the availability of open-source, multi-modal EEG emotion datasets remains limited, with DEAP and SEED being some of the few prominent resources available.
To address limitations in emotion recognition research, this study introduces the MEEG dataset, a multi-modal EEG emotion dataset inspired by the DEAP format but enhanced with music-induced emotional states for greater validity. Additionally, the LGGNet framework has been refined by integrating an attention mechanism and utilizing DGNN, significantly improving classification accuracy on the MEEG dataset. This approach facilitates a deeper exploration of interconnections between various brain functions through dynamic graph analysis.
The contributions of this study are threefold and can be summarized as follows:
  1. 1.
    We establish the MEEG dataset, a robust multi-modal EEG emotion dataset similar to DEAP but enhanced with music stimuli to effectively induce emotional states. This enhancement ensures higher ACC and F1 scores compared to DEAP using the same models.
  2. 2.
    We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain. By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion analysis tasks.
  3. 3.
    We compare the performance of the proposed method with CNN, TCN, and GNN-based SOTA methods on the MEEG dataset. Extensive ablation experiments are conducted to better understand AT-DGNN.
II Related Work
Figure 1: Structure of AT-DGNN. AT-DGNN model consists of two core modules: a feature extraction module and a dynamic graph neural network learning module. The feature extraction module (a) includes a temporal learner, multi-head attention mechanism, and a temporal convolution module. These components effectively harness local features of EEG signals through a sliding window technique, thereby enhancing the model’s ability to dynamically extract complex temporal patterns in EEG signals. In the graph learning module (b), the model first employs local filtering layers to segment and filter features from specific brain regions. Subsequently, it utilizes stacked dynamic graph convolutions to learn the intricate interactions between different brain regions, further enhancing AT-DGNN’s capability to integrate spatiotemporal features.
II-A Multi-head Attention Mechanism
EEG data is inherently complex, characterized by temporal dependencies and discrete features. Utilizing multiple attention heads allows us to capture a diverse range of temporal patterns and dependencies, effectively identifying key temporal features. In this module, the attention block consists of a multi-head attention (MHA) layer with several self-attention heads.
Each self-attention head comprises three main components: queries ($Q$), keys ($K$), and values ($V$). These components interact to generate attention scores that determine the weighted values.The procedure begins with the normalization of the input $X_{w}$ through a LayerNorm layer, a critical step for stabilizing the learning process:
$q_{ht}=k_{ht}=v_{ht}=W^{Q}\text{LN}(X_{w,t})$ (1)
This formulation demonstrates how queries, keys, and values are generated from the normalized data at each position $t$ for each head $h$, thus aiding in the computation of attention scores.
The context vector for each head is computed as a weighted sum of the values, where weights are determined by softmax-normalized alignment scores. These scores are derived using the scaled dot-product attention mechanism:
$e_{htt^{\prime}}=\frac{(q_{ht})^{T}k_{ht^{\prime}}}{\sqrt{d_{H}}}$ (2)
Ultimately, the outputs from all the heads are concatenated and subjected to a linear transformation. This consolidated output of the MHA layer is then merged with the initial input via a residual connection and normalized, leading to an optimized attention output from the layer.
II-B Graph Neural Networks
Graph neural networks (GNNs) represent a significant advancement in the domain of neural networks, designed explicitly to process graph-structured data. Unlike CNNs, GNNs excel in capturing the intricate relationships and dependencies among nodes within a graph through processes of aggregation and propagation of information across local neighbors [17]. A typical graph is denoted as $G=(V,E)$, where $V$ symbolizes the set of nodes and $E$ the set of edges. Each node $v_{i}\in V$ and edge $e_{ij}=(v_{i},v_{j})\in E$ can be respectively associated with a node and an edge in the graph. The adjacency matrix $A$ is configured as an $n\times n$ matrix where $A_{ij}=1$ if $e_{ij}\in E$ and $A_{ij}=0$ otherwise. Node attributes are represented by $X$, where $X\in\mathbb{R}^{n\times d}$, with each $x_{i}\in\mathbb{R}^{d}$ denoting the feature vector of node $v_{i}$.
DGNNs extend the capabilities of GNNs to address dynamic or time-evolving graph-structured data. In DGNNs, a graph at any given time $t$ is represented as $G_{t}=(V_{t},E_{t})$, with its corresponding adjacency matrix $A_{t}$. This matrix changes dynamically as nodes and edges are added or removed over time. Node features at time $t$ are likewise dynamic, represented as $X_{t}$ with each row $x_{t,i}\in\mathbb{R}^{d}$ embodying the evolving feature set of node $v_{i}$.
The computational heart of DGNNs lies in the dynamic update rules, where node representations $h_{t,i}$ are recurrently updated based on the temporal graph structure. A prevalent approach involves the temporal graph attention mechanism, where the new node states are computed as:
$h_{t,i}^{(k+1)}=\sigma\left(\sum_{j\in\mathcal{N}(v_{i})}\alpha_{t,ij}^{(k)}W^{(k)}h_{t,j}^{(k)}+b^{(k)}\right)$ (3)
Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma$ being a non-linear activation function.
II-C Graph Neural Networks for EEG
In 2019, Song et al. [18] first introduced a novel DGCNN specifically for multichannel EEG emotion recognition, employing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination. This method proved to be superior to existing approaches. Subsequently, advancements in GNNs have been increasingly applied to emotion recognition. On the SEED dataset, Bao et al. [19] integrated a multi-layer GNN with a style-reconfigurable CNN, while Asadzadeh et al. [20] further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance. For the DEAP dataset, Zhao et al. [21] developed coherence-based node clustering and multi-pooling techniques for GCNs. Additionally, Zhang et al. [22] enhanced GCN feature extraction by using Granger causality and differential entropy to construct more informative connectivity matrices.
III Methodology
In this section, we will introduce the detailed architecture of AT-DGNN from the three aspects of data preprocessing, feature extraction and graph learning in turn, and the overall structure of AT-DGNN is shown in Fig 1.
III-A EEG Data Preprocessing
The initial sampling rate of MEEG dataset is 1000 Hz. However, since most existing EEG datasets are sampled at either 128 Hz or 200 Hz, the data are downsampled to 200 Hz. This downsampling is crucial because the original EEG data often present significant challenges, including high levels of artifact noise, low signal amplitudes, poor interference resistance, a limited frequency range, and overlapping interference signals within the EEG band. Consequently, preprocessing of the raw EEG data is necessary before feature extraction can commence. Typically, EEG signals are categorized into five distinct frequency bands: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23]. Accordingly, a band-pass filter ranging from 1 to 50 Hz is applied to the EEG data to eliminate interference from extraneous factors.
III-B Feature Extraction
III-B1 Temporal Learner
We introduce a temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) that directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length. These kernels obviate the need for manually extracted features. The kernels’ length, dictated by the EEG data’s sampling frequency ($f_{s}$) and scaling coefficients ($\alpha^{i}$), is given by:
$S_{T}^{i}=(1,\alpha^{i}\cdot f_{s}),\quad i\in[1,2,3]$ (4)
EEG data processed through these layers yield dynamic time-frequency representations. An average pooling (AvgPool) layer, acting as a window function, calculates the averaged power across shorter segments. Applying logarithmic transformations and squaring functions further refines the feature extraction, as described by Schirrmeister et al. [10]. The output from each layer $i$, denoted as $Z^{i}_{\text{temp}}$, is formulated as:
$Z^{i}_{\text{temp}}=\Phi_{\text{log}}(\text{AvgPool}(\Phi_{\text{square}}(\mathcal{F}_{\text{Conv1-D}}(X_{i},S^{i}_{T}))))$ (5)
Here, $S^{i}_{T}$ indicates the kernel size, $\Phi_{\text{log}}$ is the logarithmic activation function, $\Phi_{\text{square}}$ represents the square function, and $\mathcal{F}_{\text{Conv1-D}}$ signifies the 1D convolution operation.
The outputs across all kernels are concatenated along the feature dimension to yield the final output of the temporal learner layer $Z_{T}$:
$Z_{T}=f_{\text{bn}}(Z^{1}_{\text{temp}},Z^{2}_{\text{temp}},Z^{3}_{\text{temp}})$ (6)
where $f_{\text{bn}}$ denotes the batch normalization operation.
III-B2 Sliding Window Segmentation
After temporal learner, we utilize a sliding window approach to segment the time series into multiple windows, which effectively extracts local features of EEG signals and augments the data. To mitigate the computational overhead introduced by the sliding window step, we employ the convolution-based sliding window method proposed by Schirrmeister et al[24]. This method integrates the sliding window approach with convolutional layers, allowing convolution operations to be performed only once across all windows, thereby reducing computation and inference time through parallel processing.
We employ a sliding window with a length of $W$ and a stride of $S$ to segment the time series $X$ into multiple windows $X_{w}\in\mathbb{R}^{B\times C\times W}$, where $w=1,\ldots,n$ denotes the window index and $n$ is the total number of windows. Each window $X_{w}$ is processed sequentially through subsequent attention and temporal convolution blocks. The formula for calculating the number of windows $n$ is:
$n=\left\lfloor\frac{\text{length}-\text{window\_size}}{\text{stride}}\right\rfloor+1$ (7)
where length is the length of the input sequence, window_size is the length of each window, and stride is the step size.
III-B3 Multi-head Attention Module
Each subsequence $X_{w}$ derived from sliding window segmentation is first standardized using a layer normalization (LayerNorm) layer. For the normalized input $X_{w}^{\text{norm}}$, the MHA mechanism computes the $q_{ht}$, $k_{ht}$, and $v_{ht}$ as linear transformations:
$\begin{cases}q_{ht}=W^{Q}\text{LN}(X_{w,t}),\\ k_{ht}=W^{K}\text{LN}(X_{w,t}),\\ v_{ht}=W^{V}\text{LN}(X_{w,t}),\end{cases}$ (8)
where LN denotes layer normalization. The matrices $W^{Q}$, $W^{K}$, and $W^{V}$ belong to $\mathbb{R}^{d\times d_{H}}$, where $d_{H}$ represents the dimension of each attention head. The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, using weights $\alpha_{htt^{\prime}}$ derived from the scaled dot-product attention mechanism.
The context vectors from all heads are concatenated and linearly transformed to yield the final output of the MHA layer:
$X_{w}^{\prime}=W^{O}[C_{1},\ldots,C_{H}]+X_{w},$ (9)
where $W^{O}\in\mathbb{R}^{d_{H}\times d}$ is the output projection matrix. This procedure effectively projects the combined outputs back to the original input dimension. Subsequently, $X_{w}^{\prime}$ is integrated with the original subsequence $X_{w}$ through a residual connection and normalized once more. This advanced MHA framework significantly enhances the model’s capacity to discern intricate temporal patterns and dependencies within EEG signals, thereby augmenting the precision of decoding activities.
III-B4 Temporal Convolution
Following the MHA mechanism, the output, denoted as attn_output, is normalized with a LayerNorm layer and merged with the sliding window data through a residual connection to generate the EEG data $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ is the number of EEG channels, and $l$ represents the sample length in the temporal dimension.
We employ multiple temporal convolutional kernels operating in parallel to extract dynamic temporal features. Each kernel’s output, $Z^{k}_{\text{tcn}}\in\mathbb{R}^{t\times c\times f_{k}}$, where $t$ denotes the total number of kernels and $f_{k}$ the feature length, acts as a digital filter capturing signals across various frequency bands. The addition of batch normalization ensures training stability, and ReLU activation introduces non-linearity.
The outputs from different kernels are concatenated along the feature dimension, forming a multi-scale output $Z_{i}^{\text{MS}}\in\mathbb{R}^{t\times c\times f_{k}}$:
$Z_{i}^{\text{MS}}=[Z^{1}_{\text{tcn}},\ldots,Z^{K}_{\text{tcn}}]$ (10)
This integration of the temporal convolution with the MHA mechanism captures local temporal dependencies effectively, enriching the feature representations and enhancing their discriminative capabilities.
III-B5 Feature Fusion
After computing the outputs of the temporal learner for each window, denoted as $Z^{w}_{\text{tcn}}\in\mathbb{R}^{B\times C\times T_{w}}$, where $B$ is the batch size, $C$ represents the channel count, $T_{w}$ is the window’s temporal length, and $w$ indexes the window, these outputs are assembled into a four-dimensional tensor:
$Z_{\text{stacked}}=[Z^{1}_{\text{tcn}},Z^{2}_{\text{tcn}},\ldots,Z^{W}_{\text{tcn}}]\in\mathbb{R}^{B\times C\times W\times T_{w}}$ (11)
This tensor is subsequently rearranged and flattened to conform to the input requirements of a convolution layer, resulting in dimensions $(B,32,-1)$, where $32$ refers to the fixed channel count and $-1$ represents the flattened dimensions. A fusion convolution layer, utilizing a kernel size of 3 and a stride of 1, integrates the outputs across all windows. The output, $Z_{\text{fused}}\in\mathbb{R}^{B\times 32\times L}$, where $L$ is the combined length, processes the temporal features to capture global contextual information effectively. This approach allows for a more discriminative representation of features in EEG signals, which is crucial for constructing nodes in GNNs.
III-C Graph Learning
III-C1 Graph Filtering Layer
For the extracted features, we adopt the method described by Ding et al.[16], which involves defining functional areas and performing local filtering. EEG data is divided into three functional areas: general region ($G_{g}$), frontal region ($G_{f}$), and hemispheric region ($G_{h}$). Each electrode is treated as a node, and the learned dynamic temporal representations are considered as node attributes. According to the 10-20 system, electrodes are grouped based on their scalp locations, as Fig 2.
Figure 2: Local global graph definitions [16]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
In the preprocessing step, EEG channels are systematically reordered within predefined groups to ensure adjacency of channels within each local graph. This reordering facilitates the subsequent application of localized graph operations. The reordering function is mathematically defined as:
$Z_{i}^{\text{reorder}}=F_{\text{reorder}}(Z_{i}^{\text{fuse}})$ (12)
Following the reordering, a graph filtering layer is employed to aggregate the transformed representations and learn patterns indicative of local brain activity. The local adjacency matrix, denoted as $A_{\text{local}}$, is initialized as a fully connected matrix where all elements are set to 1. The local graph filtering operation then applies a trainable filter matrix $W_{\text{local}}$ and a bias vector $b_{\text{local}}$ to the reordered channel representations:
$Z_{i}^{\text{filtered}}=\text{ReLU}(W_{\text{local}}\circ Z_{i}^{\text{reorder}}-b_{\text{local}})$ (13)
The output of the filtering process $Z_{i}^{\text{local}}$ within each graph is subsequently aggregated to form localized feature representations. This aggregation is performed using the defined aggregation function $F_{\text{aggregate}}$, resulting in the following vector representation for each local graph:
$Z_{i}^{\text{local}}=F_{\text{aggregate}}(Z_{i}^{\text{filtered}})=\begin{bmatrix}h_{1}^{\text{local}}\\ \vdots\\ h_{R}^{\text{local}}\end{bmatrix}$ (14)
III-C2 Stacked DGNN
Algorithm 1 AT-DGNN
1:Input: EEG data $X_{i}\in\mathbb{R}^{E\times T}$, ground truth label $y$; graph definitions $G_{g},G_{f},G_{h}$; global adjacency matrix $W_{\text{global}}$
2:Output: pred, the prediction of AT-DGNN
3:Initialization;
4$\triangleright$ Feature Extraction
5:for $i\leftarrow 1$ to $3$ do
6:     get $i$th temporal kernel size by Eq. 4;
7:     get $Z^{i}_{\text{temp}}$ by Eq. 5 using $X_{i}$ as input;
8:end for
9:get $Z_{T}$ by Eq. 6;
10:get the window size $n$ and segment the data by Eq. 7
11:get the output $X_{w}^{\prime}$ by Eq. 9 after MHA;
12:concatenate the feature dimension by Eq. 10
13:do feature fusion by Eq. 11 to get $Z_{\text{fused}}$;
14$\triangleright$ Graph Learning
15:do graph filtering and aggregation on each node by Eq. 12 - 14, get the output $Z_{i}^{\text{local}}$.
16:for $i\leftarrow 1$ to $3$ do
17:     get DGNN output by Eq. 17, Eq. 18 and Eq. 19;
18:end for
19:Get pred by Eq. 20;
20:Return pred
We propose a novel stacked DGNN architecture engineered to discern complex structural relationships within graphs. Each layer of this model adaptively recalculates the adjacency matrix using input features, facilitating the effective capture of hierarchical feature variations across multiple levels.
The architecture incorporates multiple GNN layers. In each layer, features are grouped and aggregated into matrices, with each group containing $k$ channels, resulting in an aggregated feature matrix $Z_{\text{agg}}$. The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$, defined as:
$S=Z_{\text{agg}}\cdot Z_{\text{agg}}^{T}$ (15)
To preserve node connectivity within the graph, self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is subsequently normalized as follows:
$A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (16)
During forward propagation, the initial layer performs graph convolution using the normalized adjacency matrix $\tilde{A}^{(1)}$, weight matrix $W^{(1)}$, and bias vector $b^{(1)}$. The output from this layer is given by:
$H^{(1)}=\text{ReLU}(\tilde{A}^{(1)}\cdot Z_{\text{agg}}\cdot W^{(1)}+b^{(1)})$ (17)
Subsequent layers $l=2,\ldots,L-1$ process the output of the previous layer $H^{(l-1)}$ with corresponding layer-specific parameters $\tilde{A}^{(l)}$, $W^{(l)}$, and $b^{(l)}$, to produce:
$H^{(l)}=\text{ReLU}(\tilde{A}^{(l)}\cdot H^{(l-1)}\cdot W^{(l)}+b^{(l)})$ (18)
The final output $H^{(L)}$ from the last hidden layer $H^{(L-1)}$ utilizes the parameters of the ultimate layer $\tilde{A}^{(L)}$, $W^{(L)}$, and $b^{(L)}$:
$H^{(L)}=\text{ReLU}(\tilde{A}^{(L)}\cdot H^{(L-1)}\cdot W^{(L)}+b^{(L)})$ (19)
The final network output is computed through a sequence of operations including batch normalization, dropout, and softmax activation, structured as:
$\text{Output}=\Phi_{\text{softmax}}(\mathcal{F}_{\text{dropout}}(\Gamma(\mathcal{F}_{\text{bn}}(H^{(L)}))))$ (20)
where $\Gamma(\cdot)$ denotes the flattening operation, ensuring that the network output is appropriately structured for subsequent processing or analysis.
This structured approach not only ensures detailed feature extraction at various levels but also enhances the model’s ability to learn and generalize from complex graph structures. Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
IV Experiments
TABLE I: The ACC and F1 scores (mean+std) on the MEEG dataset (trial-wise $10$-fold cross-validation)
Method Arousal Valence Arousal Valence
ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%)
EEGNet 73.40 17.82 73.03 18.31 70.00 18.06 68.52 18.95
DeepConvNet 74.99 19.91 72.04 22.86 72.38 18.87 73.03 19.27
ShallowConvNet 78.63 14.68 78.61 15.25 77.21 17.42 76.96 17.74
TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08
EEGTCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43
TCNet-Fusion 74.88 17.62 75.09 17.80 73.63 17.56 72.95 18.09
ATCNet 82.01 14.06 79.17 17.57 73.63 17.56 72.95 18.09
DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44
LGGNet-Fro 80.90 13.06 80.76 13.26 82.34 12.84 81.66 13.91
LGGNet-Gen 80.86 12.88 81.47 12.24 82.37 13.20 82.47 13.22
LGGNet-Hem 81.08 13.43 80.47 14.31 82.53 12.71 81.58 14.13
AT-DGNN-Fro 83.06 11.54 83.88 10.89 84.88 10.39 84.92 10.56
AT-DGNN-Gen 82.63 12.12 83.97 10.67 85.31 9.88 85.89 9.66
AT-DGNN-Hem 82.34 11.83 82.12 12.40 84.94 10.16 85.43 9.79
IV-A MEEG Dataset
In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments; however, participants’ responses to these stimuli may be influenced by their cultural backgrounds. Considering the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [25], we opted to induce emotions in participants using music in the MEEG dataset. We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral. To minimize emotional carryover, a 15-second pause was introduced between clips [26]. The chosen music pieces included lesser-known works of Western classical music by composers such as Shostakovich and Tchaikovsky. The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music. Each participant signed an informed consent form prior to the study. This experimental design aimed to minimize the influence of social and cultural backgrounds on emotional responses, ensuring that changes in emotions were solely due to musical stimuli, thereby enhancing the accuracy of the experimental results. EEG data were collected using a 32-channel BCI device NeuSen.W32 from Neuracle Tech, employing the same electrode channels as in the DEAP dataset [7]. The data were sampled at a frequency of 1000 Hz. The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.
IV-B Experiment Settings
To rigorously evaluate the model’s performance, a nested cross-validation approach is employed, featuring a trial-wise $10$-fold cross-validation in the outer loop and a $4$-fold cross-validation in the inner loop, as suggested by Varma [27]. This stratified sampling technique ensures robustness and generalization of the model by assessing its accuracy and reliability across diverse samples.
Additionally, a two-stage training strategy within the inner loop optimizes the utilization of training data. Initially, the best model identified from the $k$-fold cross-validation is saved as a preliminary candidate. This model is then refined using combined training data from the $k$ folds, fine-tuned at a reduced learning rate to prevent overfitting, and further trained for up to 20 epochs or until it achieves a training accuracy of 100%, ensuring precise calibration. Importantly, test data is excluded from the training process to maintain the integrity of the evaluation.
The integration of two-stage training with $10$-fold cross-validation minimizes variability in assessment results, providing a comprehensive and reliable evaluation framework suitable for diverse research and application domains.
IV-C Implement Details
The model was implemented using PyTorch [28] library, and the code can be found at https://github.com/xmh1011/AT-DGNN.
Cross-entropy loss was chosen as the objective function to guide the training process. The training was divided into two stages, with the first stage capped at 200 epochs and the second at 20 epochs. The batch size was set to 64. To reduce training time and prevent overfitting, early stopping was implemented. For the attention module, the window size was set to half the data frequency. The kernel sizes for the time learner were set to 100, 50, and 25. The hidden layer count for the DGNN was established at 32. Training was optimized using the Adam optimizer, starting with an initial learning rate of $31e-$ 3, which was reduced by a factor of 10 during the second stage. For more details, please refer to the open-source GitHub repository for AT-DGNN.
V RESULTS AND DISCUSSION
In this section, we compare the average accuracy and F1 score of AT-DGNN on the MEEG dataset with CNN, TCN, and GNN-based SOTA methods in the BCI domain. The CNN-based methods include: EEGNet [9], TSception [29], DeepConvNet and ShallowConvNet [10]. The TCN-based methods include: EEGTCNet [11], TCNet-Fusion [12] and ATCNet [30]. The GNN-based methods include: DGCNN [18] and LGGNet [16]. Due to the smaller size of the CNN-based models, the learning rate and the number of training epochs are reduced to avoid overfitting. For the other models, the parameters recommended by the authors are used. Additionally, ablation studies are conducted to reveal the contribution of each component within the AT-DGNN architecture.
V-A Emotion Recognition
Table I lists the results of our proposed method compared to other SOTA methods on the MEEG dataset. The AT-DGNN series significantly outperforms the LGGNet series across both arousal and valence dimensions. Specifically, the AT-DGNN-Gen model achieve the highest F1 score in the valence dimension at 85.89%, which represents an improvement of 3.42% over the highest score of the LGGNet series (82.47%). In the arousal dimension, AT-DGNN-Fro leads with an accuracy of 84.88%, outperforming the best LGGNet result by 2.54%. The AT-DGNN models also demonstrate greater consistency with the smallest standard deviations, indicating robustness in the challenging and class-imbalanced MEEG dataset.
Compared to traditional CNN-based methods and other machine learning techniques, the AT-DGNN series show substantial enhancements. Improvements in F1 scores are up to 17.61% higher, and accuracy gains are as much as 9.66% better, highlighting the effectiveness of the DGNN approach. These improvements underscore the capability of AT-DGNN models to better capture complex patterns and dependencies in graph-structured data, thus providing more accurate and reliable performance across varied testing conditions.
The above methods all achieve relatively high accuracy rates on the MEEG dataset, with CNN-based models reaching around 75% accuracy, or even higher. In contrast, GNN-based models have accuracy rates above 80%. According to a study by Huang et al. [31] in 2023, the same models achieve about 60% accuracy on the DEEP dataset, which is significantly lower than on the MEEG dataset. This also demonstrates that music, compared to video, is more effective in evoking emotions in subjects.
V-B Ablation Study
TABLE II: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
81.62 -1.01 82.16 -1.81
78.41 -7.16 79.99 -3.98
80.52 -2.11 81.22 -2.75
82.63 - 83.97 -
TABLE III: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
82.20 -3.17 81.46 -4.43
78.21 -7.68 78.93 -6.96
81.25 -4.12 81.10 -4.27
85.37 - 85.89 -
✓: Keep the component.
SW: Sliding window segmentation.
MHA: Multi-head attention mechanism.
D: Dynamic graph neural network.
Changes: Compared with the baseline AT-DGNN-Gen model.
We conduct ablation studies to evaluate the contributions of different components of the AT-DGNN-Gen model on the arousal classification task using MEEG data. Table II and III presents the performance of the model with various combinations of its key components: sliding window segmentation (SW), multi-head attention mechanism (MHA), and dynamic graph neural network (D). The baseline model incorporates all three components.
Removing any single component results in a notable performance drop. For the A label, excluding SW, MHA, or D reduces the ACC by 1.01%, 7.16%, and 2.11%, respectively, and the F1 score by 1.81%, 3.98%, and 2.75%, respectively. For the V label, excluding SW, MHA, or D reduces the ACC by 3.17%, 7.68%, and 4.12%, respectively, and the F1 score by 4.43%, 6.96%, and 4.27%, respectively. These findings highlight the importance of each component in enhancing the model’s ability to capture and utilize complex data relationships, validating the architectural design of AT-DGNN.
VI Conclusion
This study introduces the MEEG dataset and the AT-DGNN framework as significant advancements in EEG-based emotion recognition. The MEEG dataset, which uses music to induce emotional states, provides a valuable resource for studying brain responses to emotional stimuli. The AT-DGNN framework, integrating attention mechanisms with DGNN, excels in capturing complex temporal brain dynamics, significantly improving emotion recognition accuracy compared to other SOTA methods. These innovations advance brain-computer interface technologies and open new avenues for exploring the relationship between music, emotion, and brain activity. The superior performance of AT-DGNN, validated against traditional models, underscores its potential for enhancing emotion analysis.

### v2

@@
-MEEG and AT-DGNN: Advancing EEG Emotion Recognition with Music and Graph Learning
+MEEG and AT-DGNN: Improving EEG Emotion Recognition with Music Introducing and Graph-based Learning
 Abstract
-Recent advances in neuroscience have elucidated the crucial role of coordinated brain region activities during cognitive tasks.
-To explore this complexity, we introduce the MEEG dataset—a comprehensive multi-modal music-induced electroencephalogram (EEG) dataset and the Attention-based Temporal learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition.
-The MEEG dataset captures a wide range of emotional responses to music, enabling an in-depth analysis of brainwave patterns in musical contexts.
-The AT-DGNN combines an attention-based temporal learner with a dynamic graph neural network (DGNN) to accurately model the local and global graph dynamics of EEG data across varying brain network topologies.
-Our evaluations show that AT-DGNN achieves superior performance, with accuracies (ACC) of 83.06% in arousal and 85.31% in valence, outperforming state-of-the-art (SOTA) methods on the MEEG dataset.
-Comparative analyses with traditional datasets like DEAP highlight the effectiveness of our approach and underscore the potential of music as a powerful medium for emotion induction.
-This study not only advances our understanding of the brain’s emotional processing but also enhances the accurcy of emotion recognition technologies in brain-computer interfaces (BCI), leveraging both graph-based learning and the emotional impact of music.
-The source code and dataset are available at https://github.com/xmh1011/AT-DGNN.
-Index Terms—Electroencephalogram (EEG), Emotion recognition, Dynamic graph neural networks (DGNNs).
+Neuropsychological research highlights the essential role of coordinated activities across brain regions during cognitive tasks.
+This study expands the existing EEG datasets by constructing the MEEG dataset, a multi-modal compilation of music-induced electroencephalogram (EEG) recordings, and further investigates the brain network topology during emotional responses to music.
+The MEEG dataset, capturing emotional responses to various musical stimuli across different valence and arousal levels, enables an in-depth analysis of brainwave patterns within musical contexts.
+We introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG emotion recognition.
+By integrating an attention mechanism with a dynamic graph neural network (DGNN), the AT-DGNN model captures complex local and global EEG dynamics, demonstrating superior performance with accuracy of 83.74% in arousal and 86.01% in valence, outperforming current state-of-the-art (SOTA) methods.
+Comparative analyses with traditional datasets like DEAP underscore our model’s effectiveness, highlighting the potential of music as a potent emotional stimulus.
+This study advances graph-based learning techniques in brain-computer interfaces (BCI), significantly enhancing the precision of EEG-based emotion recognition and deepening our understanding of cognitive functions in various brain regions.
+The source code and dataset are accessible at https://github.com/xmh1011/AT-DGNN.
+Index Terms—electroencephalogram, emotion recognition, dynamic graph neural networks.
 I Introduction
@@
 In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2].
-These models conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
-BCI technology, which decodes brain signals into commands for external devices, predominantly employs EEG signals [3].
-This technology processes EEG signals through stages of preprocessing, feature extraction, and classification, thereby translating thoughts into actionable commands and enhancing human-machine interaction [4].
-EEG signals, reflecting cortical voltage fluctuations, are crucial for emotion recognition due to their objectivity, high temporal resolution, affordability, and rapid data acquisition [5].
-EEG signals have long been pivotal in neuroscience research [6].
+The model conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
+BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-machine interaction [3, 4].
+This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands.
+Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [5, 6].
 In 2011, Koelstra et al.
-established the DEAP emotion dataset by collecting physiological signals associated with emotional responses [7].
-Subsequently, in 2015, Zheng et al.
-developed the SEED dataset, providing well-annotated EEG signals from multiple subjects [8].
-Both datasets have been instrumental in facilitating the development and evaluation of computational models for emotion analysis.
-Over recent decades, diverse machine learning and signal processing techniques have been utilized for emotion recognition from physiological signals.
-A significant advancement came in 2018 when Lawhern et al.
-introduced EEGNet, a compact convolutional neural network (CNN) designed for efficient EEG signal processing [9].
-Further advancements were made by Schirrmeister et al., who enhanced EEG decoding using deep learning with CNNs and provided sophisticated visualization tools to better comprehend brain activities [10].
-Temporal convolutional networks (TCNs) have also become a formidable alternative for real-time BCI applications.
-In 2020, Ingolfsson et al.
-developed EEG-TCNet, demonstrating high accuracy in motor imagery tasks and suggesting that TCNs could surpass traditional methods in specific BCI applications [11].
-This was supported by Musallam et al., who highlighted the versatility and efficiency of TCNs in complex BCI tasks through the fusion of TCN for motor imagery classification [12].
-All the aforementioned research treated EEG signals as two-dimensional time-series data, with the dimensions being channels and time [13].
-Channels refer to EEG electrodes, which are strategically placed on the scalp to record brain activity from various functional areas according to the 10-20 system established by the International Federation of Clinical Neurophysiology [3].
-In recent years, EEG data has been treated as a graph where electrodes are positioned in a two-dimensional layout that reflects their physical placement on the scalp [14].
-In this context, the raw data or extracted features from each electrode introduce a third dimension to this spatial map [15].
+established the DEAP dataset, which consists of physiological signals, such as EEG, elicited by emotional responses to musical videos [7].
+Subsequently, Zheng et al.
+introduced the SEED dataset in 2015, comprising well-annotated EEG signals induced by cinematic stimuli, with data collected from multiple subjects [8].
+These datasets have significantly contributed to the advancement of computational models for emotion analysis.
+Over the past decades, a variety of machine learning and signal processing techniques have been applied to emotion recognition based on physiological data.
+Recent advances in deep learning have significantly impacted EEG signal processing.
+In 2018, Lawhern et al.
+proposed EEGNet, a compact convolutional neural network (CNN) tailored for EEG analysis [9].
+Subsequent developments by Schirrmeister et al.
+further refined EEG decoding capabilities using deep CNN architectures, introducing sophisticated visualization tools [10].
+The adoption of temporal convolutional networks (TCNs) has introduced robust alternatives for real-time BCI applications.
+Notably, in 2020, Ingolfsson et al.
+developed EEG-TCNet, which demonstrated superior accuracy in motor imagery tasks, suggesting that TCNs could outperform traditional methods in specific BCI contexts [11].
+This assertion was corroborated by Musallam et al., who demonstrated the versatility and efficiency of TCNs in complex BCI tasks, particularly through the integration of TCNs in motor imagery classification [12].
+All the aforementioned research treated EEG signals primarily as two-dimensional time-series data, with dimensions corresponding to channels and time [13].
+Here, channels denote EEG electrodes, positioned on the scalp using the 10-20 system, as prescribed by the International Federation of Clinical Neurophysiology, to capture neural activity from distinct brain regions [3].
+More recently, a paradigm shift has occurred with EEG data now being conceptualized as graphs, representing the electrodes’ two-dimensional arrangement on the scalp [14].
+In this model, raw data or extracted features from each electrode augment this spatial framework by adding a third dimension [15].
 In 2023, Ding et al.
-[16] introduced the Local-Global-Graph network (LGGNet), which processed EEG data in an image-like format and has shown enhanced classification performance, although it was primarily designed for multi-task classification and didn’t significantly improve emotion recognition accuracy compared to other models.
-In terms of EEG datasets, the availability of open-source, multi-modal EEG emotion datasets remains limited, with DEAP and SEED being some of the few prominent resources available.
-To address limitations in emotion recognition research, this study introduces the MEEG dataset, a multi-modal EEG emotion dataset inspired by the DEAP format but enhanced with music-induced emotional states for greater validity.
-Additionally, the LGGNet framework has been refined by integrating an attention mechanism and utilizing DGNN, significantly improving classification accuracy on the MEEG dataset.
-This approach facilitates a deeper exploration of interconnections between various brain functions through dynamic graph analysis.
-The contributions of this study are threefold and can be summarized as follows:
+introduced the Local-Global-Graph network (LGGNet) to process EEG data in an image-like format, improving classification performance [16].
+However, LGGNet’s single-layer graph neural network (GNN) struggles to capture the dynamic features of EEG signals.
+Additionally, popular EEG datasets like DEAP and SEED primarily use video stimuli, limiting the study of the relationship between music, emotions, and EEG signals.
+In contrast, EEG datasets and research using other types of stimuli are relatively scarce.
+To address shortcomings in emotion recognition research, we create the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format with music-induced emotional states to enhance ecological validity.
+We improve the LGGNet architecture by integrating an attention mechanism and using a multi-layer dynamic graph neural network, significantly boosting classification accuracy on the MEEG dataset.
+This innovative approach allows for a more nuanced examination of brain function interconnections via graph-based learning.
+The contributions of our study are threefold and can be summarized as follows:
   1. 1.
-    We establish the MEEG dataset, a robust multi-modal EEG emotion dataset similar to DEAP but enhanced with music stimuli to effectively induce emotional states.
-This enhancement ensures higher ACC and F1 scores compared to DEAP using the same models.
+    We establish the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, enhanced with various types of music to effectively induce emotional states.
+Experimental results show that the MEEG dataset outperforms the video-stimuli-based DEAP dataset in emotion induction, effectively enhancing model ACC and F1 scores.
   2. 2.
     We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain.
-By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion analysis tasks.
+By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion recognition.
   3. 3.
@@
 Figure 1: Structure of AT-DGNN.
-AT-DGNN model consists of two core modules: a feature extraction module and a dynamic graph neural network learning module.
-The feature extraction module (a) includes a temporal learner, multi-head attention mechanism, and a temporal convolution module.
-These components effectively harness local features of EEG signals through a sliding window technique, thereby enhancing the model’s ability to dynamically extract complex temporal patterns in EEG signals.
-In the graph learning module (b), the model first employs local filtering layers to segment and filter features from specific brain regions.
-Subsequently, it utilizes stacked dynamic graph convolutions to learn the intricate interactions between different brain regions, further enhancing AT-DGNN’s capability to integrate spatiotemporal features.
+The AT-DGNN model comprises two core modules: a feature extraction module (a) and a dynamic graph neural network learning module (b).
+The feature extraction module consists of a temporal learner, a multi-head attention mechanism, and a temporal convolution module.
+These components effectively leverage local features of EEG signals through a sliding window technique, thereby enhancing the model’s capacity to dynamically extract complex temporal patterns in EEG signals.
+In the graph-based learning module, the model initially employs local filtering layers to segment and filter features from specific brain regions.
+Subsequently, the architecture employs three layers of stacked dynamic graph convolutions to capture complex interactions among different brain regions.
+This structure enhances the AT-DGNN’s capacity for integrating temporal features effectively.
 II-A Multi-head Attention Mechanism
@@
 In this module, the attention block consists of a multi-head attention (MHA) layer with several self-attention heads.
-Each self-attention head comprises three main components: queries ($Q$), keys ($K$), and values ($V$).
-These components interact to generate attention scores that determine the weighted values.The procedure begins with the normalization of the input $X_{w}$ through a LayerNorm layer, a critical step for stabilizing the learning process:
-$q_{ht}=k_{ht}=v_{ht}=W^{Q}\text{LN}(X_{w,t})$ (1)
-This formulation demonstrates how queries, keys, and values are generated from the normalized data at each position $t$ for each head $h$, thus aiding in the computation of attention scores.
-The context vector for each head is computed as a weighted sum of the values, where weights are determined by softmax-normalized alignment scores.
-These scores are derived using the scaled dot-product attention mechanism:
+Each self-attention head comprises three fundamental components: queries $Q$, keys $K$, and values $V$.
+These elements facilitate the computation of attention scores that influence the weighting of the values.
+The process begins with normalizing the input $X_{w}$ via a layer normalization (LayerNorm):
+$\begin{cases}q_{ht}=W^{Q}\textit{LayerNorm}(X_{w,t})\\ k_{ht}=W^{K}\textit{LayerNorm}(X_{w,t})\\ v_{ht}=W^{V}\textit{LayerNorm}(X_{w,t})\end{cases}$ (1)
+Here, $q_{ht}$, $k_{ht}$, and $v_{ht}$ denote the queries, keys, and values at time $t$ for head $h$, derived from the normalized input.
+The matrices $W^{Q}$, $W^{K}$, and $W^{V}$ belong to $\mathbb{R}^{d\times d_{H}}$, where $d_{H}$ represents the dimension of each attention head.
+This normalization facilitates the efficient computation of attention scores.
+The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, with weights $\alpha_{htt^{\prime}}$ determined by the scaled dot-product attention mechanism.
+The alignment scores $e_{htt^{\prime}}$ are calculated as:
 $e_{htt^{\prime}}=\frac{(q_{ht})^{T}k_{ht^{\prime}}}{\sqrt{d_{H}}}$ (2)
-Ultimately, the outputs from all the heads are concatenated and subjected to a linear transformation.
-This consolidated output of the MHA layer is then merged with the initial input via a residual connection and normalized, leading to an optimized attention output from the layer.
+Here, $t^{\prime}$ denotes the different time steps considered. The softmax function converts these scores into a probability distribution, reflecting the importance of each value vector:
+$c_{ht}=\sum_{t^{\prime}}\alpha_{htt^{\prime}}v_{ht^{\prime}}$ (3)
+where $\alpha_{htt^{\prime}}$ represents the normalized alignment scores, indicating the importance of each value vector $v_{ht^{\prime}}$ in forming the context vector $c_{ht}$ for time step $t$.
+This context vector $c_{ht}$ captures the most relevant information from the values, highlighting key features at each time step.
 II-B Graph Neural Networks
-Graph neural networks (GNNs) represent a significant advancement in the domain of neural networks, designed explicitly to process graph-structured data.
+GNNs represent a significant advancement in the domain of neural networks, designed explicitly to process graph-structured data.
 Unlike CNNs, GNNs excel in capturing the intricate relationships and dependencies among nodes within a graph through processes of aggregation and propagation of information across local neighbors [17].
@@
 A prevalent approach involves the temporal graph attention mechanism, where the new node states are computed as:
-$h_{t,i}^{(k+1)}=\sigma\left(\sum_{j\in\mathcal{N}(v_{i})}\alpha_{t,ij}^{(k)}W^{(k)}h_{t,j}^{(k)}+b^{(k)}\right)$ (3)
-Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma$ being a non-linear activation function.
+$\textit{h}_{t,i}^{(k+1)}=\sigma\left(\sum_{j\in\mathcal{N}(v_{i})}\alpha_{t,ij}^{(k)}W^{(k)}\textit{h}_{t,j}^{(k)}+\textit{b}^{(k)}\right)$ (4)
+Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma(\cdot)$ being a non-linear activation function.
 II-C Graph Neural Networks for EEG
@@
 Additionally, Zhang et al.
-[22] enhanced GCN feature extraction by using Granger causality and differential entropy to construct more informative connectivity matrices.
+[22] enhanced GCN feature extraction by using Granger causality analysis and differential entropy to construct more informative connectivity matrices.
 III Methodology
-In this section, we will introduce the detailed architecture of AT-DGNN from the three aspects of data preprocessing, feature extraction and graph learning in turn, and the overall structure of AT-DGNN is shown in Fig 1.
+In this section, we introduce the detailed architecture of AT-DGNN from the three aspects of data preprocessing, feature extraction and graph-based learning in turn, and the overall structure of AT-DGNN is shown in Fig.
+1.
 III-A EEG Data Preprocessing
-The initial sampling rate of MEEG dataset is 1000 Hz.
-However, since most existing EEG datasets are sampled at either 128 Hz or 200 Hz, the data are downsampled to 200 Hz.
-This downsampling is crucial because the original EEG data often present significant challenges, including high levels of artifact noise, low signal amplitudes, poor interference resistance, a limited frequency range, and overlapping interference signals within the EEG band.
-Consequently, preprocessing of the raw EEG data is necessary before feature extraction can commence.
-Typically, EEG signals are categorized into five distinct frequency bands: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23].
-Accordingly, a band-pass filter ranging from 1 to 50 Hz is applied to the EEG data to eliminate interference from extraneous factors.
+The initial sampling rate of the MEEG dataset is 1000 Hz.
+To align with standard EEG datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz.
+This preprocessing phase involves the use of a band-pass filter with a range of 1 to 50 Hz to enhance the signal quality by minimizing interference.
+Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23].
 III-B Feature Extraction
 III-B1 Temporal Learner
-We introduce a temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) that directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length.
+A temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) is used to directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length.
 These kernels obviate the need for manually extracted features.
-The kernels’ length, dictated by the EEG data’s sampling frequency ($f_{s}$) and scaling coefficients ($\alpha^{i}$), is given by:
+The $i$th-kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), is given by [24]:
 $S_{T}^{i}=(1,\alpha^{i}\cdot f_{s}),\quad i\in[1,2,3]$ (5)
@@
 An average pooling (AvgPool) layer, acting as a window function, calculates the averaged power across shorter segments.
-Applying logarithmic transformations and squaring functions further refines the feature extraction, as described by Schirrmeister et al.
-[10].
-The output from each layer $i$, denoted as $Z^{i}_{\text{temp}}$, is formulated as:
-$Z^{i}_{\text{temp}}=\Phi_{\text{log}}(\text{AvgPool}(\Phi_{\text{square}}(\mathcal{F}_{\text{Conv1-D}}(X_{i},S^{i}_{T}))))$ (5)
-Here, $S^{i}_{T}$ indicates the kernel size, $\Phi_{\text{log}}$ is the logarithmic activation function, $\Phi_{\text{square}}$ represents the square function, and $\mathcal{F}_{\text{Conv1-D}}$ signifies the 1D convolution operation.
+The logarithmic activation as described by [10] is applied to improve the performance.
+The output from each layer $i$, denoted as $Z^{i}_{\textit{temp}}$, is formulated as:
+$Z^{i}_{\textit{temp}}=\Phi_{\textit{log}}(\textit{AvgPool}(\Phi_{\textit{sqr}}(\mathcal{F}_{\textit{Conv1-D}}(X_{i},S^{i}_{T}))))$ (6)
+Here, $\Phi_{\textit{log}}(\cdot)$ is the logarithmic activation function, $\Phi_{\textit{sqr}}(\cdot)$ represents the square function, and $\mathcal{F}_{\textit{Conv1-D}}(\cdot)$ signifies the 1D convolution operation.
 The outputs across all kernels are concatenated along the feature dimension to yield the final output of the temporal learner layer $Z_{T}$:
-$Z_{T}=f_{\text{bn}}(Z^{1}_{\text{temp}},Z^{2}_{\text{temp}},Z^{3}_{\text{temp}})$ (6)
-where $f_{\text{bn}}$ denotes the batch normalization operation.
+$Z_{T}=f_{\textit{bn}}(Z^{1}_{\textit{temp}},Z^{2}_{\textit{temp}},Z^{3}_{\textit{temp}})$ (7)
+where $f_{\textit{bn}}(\cdot)$ denotes the batch normalization operation.
 III-B2 Sliding Window Segmentation
-After temporal learner, we utilize a sliding window approach to segment the time series into multiple windows, which effectively extracts local features of EEG signals and augments the data.
-To mitigate the computational overhead introduced by the sliding window step, we employ the convolution-based sliding window method proposed by Schirrmeister et al[24].
-This method integrates the sliding window approach with convolutional layers, allowing convolution operations to be performed only once across all windows, thereby reducing computation and inference time through parallel processing.
-We employ a sliding window with a length of $W$ and a stride of $S$ to segment the time series $X$ into multiple windows $X_{w}\in\mathbb{R}^{B\times C\times W}$, where $w=1,\ldots,n$ denotes the window index and $n$ is the total number of windows.
-Each window $X_{w}$ is processed sequentially through subsequent attention and temporal convolution blocks.
-The formula for calculating the number of windows $n$ is:
-$n=\left\lfloor\frac{\text{length}-\text{window\_size}}{\text{stride}}\right\rfloor+1$ (7)
-where length is the length of the input sequence, window_size is the length of each window, and stride is the step size.
+Following the temporal learner, a convolution-based sliding window technique as [25] is applied to segment the EEG time series $X$ into multiple windows.
+This method integrates sliding window segmentation with convolutional operations, significantly reducing computational overhead by allowing convolutions to be executed once across all windows.
+This approach not only enhances data augmentation but also accelerates processing through parallel computation.
+The time series $X$ is segmented into windows $X_{w}\in\mathbb{R}^{B\times C\times W}$ using a window of length $W$ and stride $S$, where $w=1,\ldots,n$ represents the window index and $n$ is the total number of windows.
+Each window $X_{w}$ is then processed by subsequent attention and temporal convolution blocks.
+The number of windows $n$ is calculated as:
+$n=\left\lfloor\frac{\textit{length}-\textit{window\_size}}{\textit{stride}}\right\rfloor+1$ (8)
+Here, length, window_size, and stride denote the total length of the time series, the length of each window, and the stride between windows, respectively.
 III-B3 Multi-head Attention Module
-Each subsequence $X_{w}$ derived from sliding window segmentation is first standardized using a layer normalization (LayerNorm) layer.
-For the normalized input $X_{w}^{\text{norm}}$, the MHA mechanism computes the $q_{ht}$, $k_{ht}$, and $v_{ht}$ as linear transformations:
-$\begin{cases}q_{ht}=W^{Q}\text{LN}(X_{w,t}),\\ k_{ht}=W^{K}\text{LN}(X_{w,t}),\\ v_{ht}=W^{V}\text{LN}(X_{w,t}),\end{cases}$ (8)
-where LN denotes layer normalization.
-The matrices $W^{Q}$, $W^{K}$, and $W^{V}$ belong to $\mathbb{R}^{d\times d_{H}}$, where $d_{H}$ represents the dimension of each attention head.
-The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, using weights $\alpha_{htt^{\prime}}$ derived from the scaled dot-product attention mechanism.
-The context vectors from all heads are concatenated and linearly transformed to yield the final output of the MHA layer:
-$X_{w}^{\prime}=W^{O}[C_{1},\ldots,C_{H}]+X_{w},$ (9)
+Each subsequence $X_{w}$ derived from sliding window segmentation is processed as described in (1).
+The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, with weights $\alpha_{htt^{\prime}}$ determined by the scaled dot-product attention mechanism as (2), and then used in (3).
+The context vectors $c_{1t},\ldots,c_{Ht}$ from all heads are concatenated and linearly transformed to yield the final output of the MHA layer:
+$\textit{X}_{w}^{\prime}=W^{O}[\textit{c}_{1t},\ldots,\textit{c}_{Ht}]+\textit{X}_{w},$ (9)
 where $W^{O}\in\mathbb{R}^{d_{H}\times d}$ is the output projection matrix.
-This procedure effectively projects the combined outputs back to the original input dimension.
+This procedure projects the combined outputs back to the original input dimension.
 Subsequently, $X_{w}^{\prime}$ is integrated with the original subsequence $X_{w}$ through a residual connection and normalized once more.
@@
 III-B4 Temporal Convolution
-Following the MHA mechanism, the output, denoted as attn_output, is normalized with a LayerNorm layer and merged with the sliding window data through a residual connection to generate the EEG data $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ is the number of EEG channels, and $l$ represents the sample length in the temporal dimension.
-We employ multiple temporal convolutional kernels operating in parallel to extract dynamic temporal features.
-Each kernel’s output, $Z^{k}_{\text{tcn}}\in\mathbb{R}^{t\times c\times f_{k}}$, where $t$ denotes the total number of kernels and $f_{k}$ the feature length, acts as a digital filter capturing signals across various frequency bands.
-The addition of batch normalization ensures training stability, and ReLU activation introduces non-linearity.
-The outputs from different kernels are concatenated along the feature dimension, forming a multi-scale output $Z_{i}^{\text{MS}}\in\mathbb{R}^{t\times c\times f_{k}}$:
-$Z_{i}^{\text{MS}}=[Z^{1}_{\text{tcn}},\ldots,Z^{K}_{\text{tcn}}]$ (10)
-This integration of the temporal convolution with the MHA mechanism captures local temporal dependencies effectively, enriching the feature representations and enhancing their discriminative capabilities.
+Following the MHA mechanism, the output $\textit{X}_{w}^{\prime}$ is normalized using LayerNorm.
+The normalized EEG data are denoted as $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ represents the number of EEG channels and $l$ represents the temporal dimension’s sample length.
+To extract dynamic temporal features, we employ a temporal convolutional block on each window of the EEG data.
+The normalized EEG data $X_{i}$ serves as the input to this block.
+The temporal convolutional block processes $X_{i}$ and generates an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ represents the window’s temporal length and $w$ indexes the window.
+Batch normalization is applied to the output to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity.
+The output $Z_{\textit{tcn}}^{w}$ is computed as follows:
+$Z_{\textit{tcn}}^{w}=\textit{ReLU}(\textit{BatchNorm}(\textit{Conv1d}(X_{i})))$ (10)
 III-B5 Feature Fusion
-After computing the outputs of the temporal learner for each window, denoted as $Z^{w}_{\text{tcn}}\in\mathbb{R}^{B\times C\times T_{w}}$, where $B$ is the batch size, $C$ represents the channel count, $T_{w}$ is the window’s temporal length, and $w$ indexes the window, these outputs are assembled into a four-dimensional tensor:
-$Z_{\text{stacked}}=[Z^{1}_{\text{tcn}},Z^{2}_{\text{tcn}},\ldots,Z^{W}_{\text{tcn}}]\in\mathbb{R}^{B\times C\times W\times T_{w}}$ (11)
-This tensor is subsequently rearranged and flattened to conform to the input requirements of a convolution layer, resulting in dimensions $(B,32,-1)$, where $32$ refers to the fixed channel count and $-1$ represents the flattened dimensions.
-A fusion convolution layer, utilizing a kernel size of 3 and a stride of 1, integrates the outputs across all windows.
-The output, $Z_{\text{fused}}\in\mathbb{R}^{B\times 32\times L}$, where $L$ is the combined length, processes the temporal features to capture global contextual information effectively.
-This approach allows for a more discriminative representation of features in EEG signals, which is crucial for constructing nodes in GNNs.
-III-C Graph Learning
+The outputs of the temporal learner for each window, denoted as $Z^{w}_{\textit{tcn}}\in\mathbb{R}^{B\times C\times T_{w}}$, are assembled into a four-dimensional tensor $Z_{\textit{stacked}}$:
+$Z_{\textit{stacked}}=[Z^{1}_{\textit{tcn}},Z^{2}_{\textit{tcn}},\ldots,Z^{W}_{\textit{tcn}}]\in\mathbb{R}^{B\times C\times W\times T_{w}}$ (11)
+This tensor is rearranged and flattened to meet the convolution layer’s input requirements, resulting in dimensions $(B,32,-1)$, where $32$ is the fixed channel count, and $-1$ represents the flattened dimensions.
+A fusion convolution layer with a kernel size of 3 and a stride of 1 integrates the outputs across all windows.
+The final output $Z_{\text{fused}}\in\mathbb{R}^{B\times 32\times L}$, where $L$ is the combined length, captures global contextual information, allowing for a more discriminative representation of features in EEG signals, which is crucial for constructing nodes in GNNs.
+III-C Graph-based Learning
 III-C1 Graph Filtering Layer
-For the extracted features, we adopt the method described by Ding et al.[16], which involves defining functional areas and performing local filtering.
-EEG data is divided into three functional areas: general region ($G_{g}$), frontal region ($G_{f}$), and hemispheric region ($G_{h}$).
-Each electrode is treated as a node, and the learned dynamic temporal representations are considered as node attributes.
-According to the 10-20 system, electrodes are grouped based on their scalp locations, as Fig 2.
+For the extracted features, we adopt the method described by [16], which involves defining functional areas and performing local filtering.
+The electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig.
+2.
+Each channel is treated as a node, and the learned dynamic temporal representations are considered as node attributes.
 Figure 2: Local global graph definitions [16]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
 In the preprocessing step, EEG channels are systematically reordered within predefined groups to ensure adjacency of channels within each local graph.
-This reordering facilitates the subsequent application of localized graph operations.
-The reordering function is mathematically defined as:
-$Z_{i}^{\text{reorder}}=F_{\text{reorder}}(Z_{i}^{\text{fuse}})$ (12)
-Following the reordering, a graph filtering layer is employed to aggregate the transformed representations and learn patterns indicative of local brain activity.
-The local adjacency matrix, denoted as $A_{\text{local}}$, is initialized as a fully connected matrix where all elements are set to 1.
-The local graph filtering operation then applies a trainable filter matrix $W_{\text{local}}$ and a bias vector $b_{\text{local}}$ to the reordered channel representations:
-$Z_{i}^{\text{filtered}}=\text{ReLU}(W_{\text{local}}\circ Z_{i}^{\text{reorder}}-b_{\text{local}})$ (13)
-The output of the filtering process $Z_{i}^{\text{local}}$ within each graph is subsequently aggregated to form localized feature representations.
-This aggregation is performed using the defined aggregation function $F_{\text{aggregate}}$, resulting in the following vector representation for each local graph:
-$Z_{i}^{\text{local}}=F_{\text{aggregate}}(Z_{i}^{\text{filtered}})=\begin{bmatrix}h_{1}^{\text{local}}\\ \vdots\\ h_{R}^{\text{local}}\end{bmatrix}$ (14)
+This reordering enhances the effectiveness of localized graph-based operations applied in subsequent stages.
+Mathematically, the reordering of channel representations, denoted by $Z_{\textit{reorder}}^{i}$, is defined by the function $F_{\textit{reorder}}(\cdot)$, which operates on the fused data from EEG channels at the $i$-th observation representations $Z_{\textit{fuse}}^{i}$.
+This relationship is expressed as:
+$Z_{\textit{reorder}}^{i}=F_{\textit{reorder}}(Z_{\textit{fuse}}^{i})$ (12)
+After reordering, a graph filtering layer is employed to aggregate the transformed representations of EEG channels using a fully connected local adjacency matrix $A_{\textit{local}}$, where all elements are set to 1.
+The filtered output $Z_{\textit{filtered}}^{i}$ is obtained by applying a trainable filter matrix $W_{\textit{local}}$ and a bias $b_{\textit{local}}$ to the reordered data, followed by the application of the $\textit{ReLU}(\cdot)$ activation function:
+$Z_{\textit{filtered}}^{i}=\textit{ReLU}(W_{\textit{local}}\circ Z_{\textit{reorder}}^{i}+b_{\textit{local}})$ (13)
+The element-wise product symbol $\circ$ captures localized interactions within the brain.
+The output of the filtering process, $Z_{\textit{local}}^{i}$, is aggregated to form localized feature representations using the aggregation function $F_{\textit{aggregate}}(\cdot)$.
+The vector representation for each local graph is:
+$Z_{\textit{local}}^{i}=F_{\textit{aggregate}}(Z_{\textit{filtered}}^{i})=\begin{bmatrix}h_{\textit{local}}^{1}\\ \vdots\\ h_{\textit{local}}^{R}\end{bmatrix}$ (14)
+where $h_{\textit{local}}^{1},\dots,h_{\textit{local}}^{R}$ are elements of the localized feature representation vector.
+These features $Z_{\textit{local}}^{i}$ are used to construct the overall feature matrix $Z_{\textit{agg}}$:
+$Z_{\textit{agg}}=\begin{bmatrix}Z_{\textit{local}}^{1}\\ Z_{\textit{local}}^{2}\\ \vdots\\ Z_{\textit{local}}^{11}\end{bmatrix}$ (15)
 III-C2 Stacked DGNN
 Algorithm 1 AT-DGNN
-1:Input: EEG data $X_{i}\in\mathbb{R}^{E\times T}$, ground truth label $y$; graph definitions $G_{g},G_{f},G_{h}$; global adjacency matrix $W_{\text{global}}$
+1:Input: EEG data $X_{i}\in\mathbb{R}^{E\times T}$, ground truth label $y$
 2:Output: pred, the prediction of AT-DGNN
@@
 5:for $i\leftarrow 1$ to $3$ do
-6:     get $i$th temporal kernel size by Eq. 4;
-7:     get $Z^{i}_{\text{temp}}$ by Eq. 5 using $X_{i}$ as input;
+6:     get $i$th temporal kernel size by (5);
+7:     get $Z^{i}_{\textit{temp}}$ by (6) using $X_{i}$ as input;
 8:end for
-9:get $Z_{T}$ by Eq. 6;
-10:get the window size $n$ and segment the data by Eq. 7
-11:get the output $X_{w}^{\prime}$ by Eq. 9 after MHA;
-12:concatenate the feature dimension by Eq. 10
-13:do feature fusion by Eq. 11 to get $Z_{\text{fused}}$;
-14$\triangleright$ Graph Learning
-15:do graph filtering and aggregation on each node by Eq. 12 - 14, get the output $Z_{i}^{\text{local}}$.
-16:for $i\leftarrow 1$ to $3$ do
-17:     get DGNN output by Eq. 17, Eq. 18 and Eq. 19;
+9:get $Z_{T}$ by (7);
+10:get the window size $n$ and segment the data by (8):
+11:for $w\leftarrow 1$ to $n$ do
+12:     get window $X_{w}\in\mathbb{R}^{B\times E\times W}$;
+13:     normalize $X_{w}$ using LayerNorm;
+14:     process $X_{w}$ with MHA to get $X_{w}^{\prime}$ by (9);
+15:     normalize $X_{w}^{\prime}$ using LayerNorm;
+16:     apply TCN on $X_{w}^{\prime}$ to get $Z_{\textit{tcn}}^{w}$ by (10);
+17:     add $Z_{\textit{tcn}}^{w}$ to $Z_{\textit{stacked}}$;
 18:end for
-19:Get pred by Eq. 20;
-20:Return pred
-We propose a novel stacked DGNN architecture engineered to discern complex structural relationships within graphs.
-Each layer of this model adaptively recalculates the adjacency matrix using input features, facilitating the effective capture of hierarchical feature variations across multiple levels.
+19:apply Fusion Conv to $Z_{\textit{stacked}}$ to get $Z_{\textit{fused}}$ by (11);
+20$\triangleright$ Graph-based Learning
+21:do graph filtering and aggregation on each node by (12) - 15 to get $Z^{i}_{\textit{local}}$;
+22:for $i\leftarrow 1$ to $3$ do
+23:     get DGNN output by (18);
+24:end for
+25:get pred by (19);
+26:Return pred
+We introduce a stacked DGNN architecture to discern complex structural relationships within graphs.
+Each layer adaptively recalculates the adjacency matrix based on input features, effectively capturing hierarchical feature variations.
 The architecture incorporates multiple GNN layers.
-In each layer, features are grouped and aggregated into matrices, with each group containing $k$ channels, resulting in an aggregated feature matrix $Z_{\text{agg}}$.
-The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$, defined as:
-$S=Z_{\text{agg}}\cdot Z_{\text{agg}}^{T}$ (15)
-To preserve node connectivity within the graph, self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is subsequently normalized as follows:
+In each layer, features are grouped and aggregated into matrices, resulting in an aggregated feature matrix $Z_{\textit{agg}}$.
+The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$:
+$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
+Self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is normalized as:
 $A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (17)
-During forward propagation, the initial layer performs graph convolution using the normalized adjacency matrix $\tilde{A}^{(1)}$, weight matrix $W^{(1)}$, and bias vector $b^{(1)}$.
-The output from this layer is given by:
-$H^{(1)}=\text{ReLU}(\tilde{A}^{(1)}\cdot Z_{\text{agg}}\cdot W^{(1)}+b^{(1)})$ (17)
-Subsequent layers $l=2,\ldots,L-1$ process the output of the previous layer $H^{(l-1)}$ with corresponding layer-specific parameters $\tilde{A}^{(l)}$, $W^{(l)}$, and $b^{(l)}$, to produce:
-$H^{(l)}=\text{ReLU}(\tilde{A}^{(l)}\cdot H^{(l-1)}\cdot W^{(l)}+b^{(l)})$ (18)
-The final output $H^{(L)}$ from the last hidden layer $H^{(L-1)}$ utilizes the parameters of the ultimate layer $\tilde{A}^{(L)}$, $W^{(L)}$, and $b^{(L)}$:
-$H^{(L)}=\text{ReLU}(\tilde{A}^{(L)}\cdot H^{(L-1)}\cdot W^{(L)}+b^{(L)})$ (19)
-The final network output is computed through a sequence of operations including batch normalization, dropout, and softmax activation, structured as:
-$\text{Output}=\Phi_{\text{softmax}}(\mathcal{F}_{\text{dropout}}(\Gamma(\mathcal{F}_{\text{bn}}(H^{(L)}))))$ (20)
+Here, $\tilde{D}$ is the diagonal degree matrix of $A$, with each diagonal element $\tilde{D}_{ii}$ being the sum of the $i$-th row of $A$.
+During forward propagation, each DGNN layer processes the inputs using a normalized adjacency matrix $\tilde{A}_{i}$, a weight matrix $W_{i}$, and a bias vector $b_{i}$ for graph convolution.
+The outputs for each layer are calculated as:
+$\begin{cases}H_{1}=\textit{ReLU}(\tilde{A}_{1}\cdot Z_{\textit{agg}}\cdot W_{1}+b_{1})\\ H_{i}=\textit{ReLU}(\tilde{A}_{i}\cdot H_{i-1}\cdot W_{i}+b_{i}),\quad i\in[2,n-1]\\ H_{n}=\textit{ReLU}(\tilde{A}_{n}\cdot H_{n-1}\cdot W_{n}+b_{n})\end{cases}$ (18)
+Here, $n$ denotes the total number of layers in the DGNN, which is set to $3$ in our study.
+The final Output is computed through a sequence of operations including batch normalization $\mathcal{F}_{\textit{bn}}(\cdot)$, dropout $\mathcal{F}_{\textit{dropout}}(\cdot)$, and softmax activation $\Phi_{\textit{softmax}}(\cdot)$, structured as:
+$\textit{Output}=\Phi_{\textit{softmax}}(\mathcal{F}_{\textit{dropout}}(\Gamma(\mathcal{F}_{\textit{bn}}(H^{(L)}))))$ (19)
 where $\Gamma(\cdot)$ denotes the flattening operation, ensuring that the network output is appropriately structured for subsequent processing or analysis.
-This structured approach not only ensures detailed feature extraction at various levels but also enhances the model’s ability to learn and generalize from complex graph structures.
 Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
 IV Experiments
-TABLE I: The ACC and F1 scores (mean+std) on the MEEG dataset (trial-wise $10$-fold cross-validation)
-Method Arousal Valence Arousal Valence
-ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%)
-EEGNet 73.40 17.82 73.03 18.31 70.00 18.06 68.52 18.95
-DeepConvNet 74.99 19.91 72.04 22.86 72.38 18.87 73.03 19.27
-ShallowConvNet 78.63 14.68 78.61 15.25 77.21 17.42 76.96 17.74
-TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08
-EEGTCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43
-TCNet-Fusion 74.88 17.62 75.09 17.80 73.63 17.56 72.95 18.09
-ATCNet 82.01 14.06 79.17 17.57 73.63 17.56 72.95 18.09
-DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44
-LGGNet-Fro 80.90 13.06 80.76 13.26 82.34 12.84 81.66 13.91
-LGGNet-Gen 80.86 12.88 81.47 12.24 82.37 13.20 82.47 13.22
-LGGNet-Hem 81.08 13.43 80.47 14.31 82.53 12.71 81.58 14.13
-AT-DGNN-Fro 83.06 11.54 83.88 10.89 84.88 10.39 84.92 10.56
-AT-DGNN-Gen 82.63 12.12 83.97 10.67 85.31 9.88 85.89 9.66
-AT-DGNN-Hem 82.34 11.83 82.12 12.40 84.94 10.16 85.43 9.79
 IV-A MEEG Dataset
-In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments; however, participants’ responses to these stimuli may be influenced by their cultural backgrounds.
-Considering the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [25], we opted to induce emotions in participants using music in the MEEG dataset.
+In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments.
+However, participants’ responses to these stimuli may be influenced by their cultural backgrounds [26].
+Considering that the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [27], we opted to induce emotions in participants using music in the MEEG dataset.
 We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral.
@@
 For more details, please refer to the open-source GitHub repository for AT-DGNN.
+TABLE I: Comparison of ACC and F1 scores (mean $\pm$ std) for the MEEG dataset under trial-wise $10$-fold cross-validation
+Method Arousal Valence Emotion
+ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%) ACC (%) F1 (%)
+EEGNet 75.47 17.63 75.54 17.83 73.66 16.82 73.33 16.94 74.57 74.44
+DeepConvNet 79.35 15.03 78.43 18.10 72.38 18.87 73.03 19.27 75.87 75.73
+ShallowConvNet 80.96 14.49 80.87 14.71 79.77 16.44 80.22 15.60 80.37 80.55
+TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08 78.79 81.24
+EEG-TCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43 73.04 64.04
+TCNet-Fusion 76.72 17.32 76.95 17.49 75.88 16.70 75.42 16.97 76.30 76,20
+ATCNet 82.01 14.06 79.17 17.57 83.19 12.09 82.08 13.41 82.60 80.63
+DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44 82.42 82.12
+LGGNet-Fro 81.85 12.57 81.58 12.73 82.34 12.84 81.66 13.91 82.11 81.62
+LGGNet-Gen 82.15 12.72 82.15 12.57 82.53 13.20 82.47 13.22 82.34 82.31
+LGGNet-Hem 81.92 12.08 81.58 12.43 82.37 12.71 81.58 14.13 82.15 81.58
+AT-DGNN-Fro 83.51 11.09 83.06 11.83 85.56 10.69 85.61 10.89 84.54 84.34
+AT-DGNN-Gen 83.74 12.03 84.56 10.88 86.01 9.40 85.68 10.06 84.88 85.12
+AT-DGNN-Hem 83.73 11.20 84.78 10.93 84.89 10.57 85.35 10.24 84.31 85.07
+  • Emotion averages the ACC and F1 scores of both the arousal and valence dimensions.
+  • The results for all models are derived from the optimal values among ten trials with distinct random seeds.
 V RESULTS AND DISCUSSION
-In this section, we compare the average accuracy and F1 score of AT-DGNN on the MEEG dataset with CNN, TCN, and GNN-based SOTA methods in the BCI domain.
+In this section, we compare the average ACC and F1 score of AT-DGNN on the MEEG dataset with CNN, TCN, and GNN-based SOTA methods in the BCI domain.
 The CNN-based methods include: EEGNet [9], TSception [24], DeepConvNet and ShallowConvNet [10].
-The TCN-based methods include: EEGTCNet [11], TCNet-Fusion [12] and ATCNet [30].
+The TCN-based methods include: EEG-TCNet [11], TCNet-Fusion [12] and ATCNet [31].
 The GNN-based methods include: DGCNN [18] and LGGNet [16].
@@
 V-A Emotion Recognition
-Table I lists the results of our proposed method compared to other SOTA methods on the MEEG dataset.
-The AT-DGNN series significantly outperforms the LGGNet series across both arousal and valence dimensions.
-Specifically, the AT-DGNN-Gen model achieve the highest F1 score in the valence dimension at 85.89%, which represents an improvement of 3.42% over the highest score of the LGGNet series (82.47%).
-In the arousal dimension, AT-DGNN-Fro leads with an accuracy of 84.88%, outperforming the best LGGNet result by 2.54%.
-The AT-DGNN models also demonstrate greater consistency with the smallest standard deviations, indicating robustness in the challenging and class-imbalanced MEEG dataset.
-Compared to traditional CNN-based methods and other machine learning techniques, the AT-DGNN series show substantial enhancements.
-Improvements in F1 scores are up to 17.61% higher, and accuracy gains are as much as 9.66% better, highlighting the effectiveness of the DGNN approach.
-These improvements underscore the capability of AT-DGNN models to better capture complex patterns and dependencies in graph-structured data, thus providing more accurate and reliable performance across varied testing conditions.
-The above methods all achieve relatively high accuracy rates on the MEEG dataset, with CNN-based models reaching around 75% accuracy, or even higher.
-In contrast, GNN-based models have accuracy rates above 80%.
-According to a study by Huang et al.
-[31] in 2023, the same models achieve about 60% accuracy on the DEEP dataset, which is significantly lower than on the MEEG dataset.
-This also demonstrates that music, compared to video, is more effective in evoking emotions in subjects.
+As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy.
+The AT-DGNN-Gen model not only achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, but also records the highest accuracy of 86.01% in the valence dimension, a substantial improvement of 3.48% over the LGGNet series.
+This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
+The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions.
+This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks.
+Furthermore, consistent with the findings of [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
+Furthermore, these results demonstrate a marked improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy surpassing previous models by up to 7.16% and 5.78%, respectively.
+This emphasizes the advanced capability of the DGNN approach in graph-structured data analysis, providing more accurate and reliable emotion recognition.
+All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%.
+On the DEAP dataset, AT-DGNN averages 60.56% accuracy across four dimensions.
+According to Huang et al.
+[32], other models, including LGGNet, also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not.
+All models perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
 V-B Ablation Study
-TABLE II: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
-SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
-81.62 -1.01 82.16 -1.81
-78.41 -7.16 79.99 -3.98
-80.52 -2.11 81.22 -2.75
-82.63 - 83.97 -
-TABLE III: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
-SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
-82.20 -3.17 81.46 -4.43
-78.21 -7.68 78.93 -6.96
-81.25 -4.12 81.10 -4.27
-85.37 - 85.89 -
+TABLE II: Ablation study on Emotion of MEEG using AT-DGNN-Gen.
+S M D ACC(%) Changes(%) F1(%) Changes(%)
+81.91 -2.97 81.81 -3.31
+3 78.31 -6.57 79.46 -5.66
+3 80.89 -3.99 81.16 -3.96
+1 81.91 -2.97 81.81 -3.31
+2 82.94 -1.94 82.52 -2.60
+4 83.17 -1.71 83.48 -1.64
+3 84.88 - 85.12 -
+  • S: Sliding window segmentation.
+  • M: Multi-head attention mechanism.
+  • D: Dynamic graph neural network.
   • ✓: Keep the component.
-SW: Sliding window segmentation.
-MHA: Multi-head attention mechanism.
-D: Dynamic graph neural network.
-Changes: Compared with the baseline AT-DGNN-Gen model.
-We conduct ablation studies to evaluate the contributions of different components of the AT-DGNN-Gen model on the arousal classification task using MEEG data.
-Table II and III presents the performance of the model with various combinations of its key components: sliding window segmentation (SW), multi-head attention mechanism (MHA), and dynamic graph neural network (D).
-The baseline model incorporates all three components.
-Removing any single component results in a notable performance drop.
-For the A label, excluding SW, MHA, or D reduces the ACC by 1.01%, 7.16%, and 2.11%, respectively, and the F1 score by 1.81%, 3.98%, and 2.75%, respectively.
-For the V label, excluding SW, MHA, or D reduces the ACC by 3.17%, 7.68%, and 4.12%, respectively, and the F1 score by 4.43%, 6.96%, and 4.27%, respectively.
-These findings highlight the importance of each component in enhancing the model’s ability to capture and utilize complex data relationships, validating the architectural design of AT-DGNN.
+  • ✗: Exclude the component.
+  • 1, 2, 3, 4: The number of stacked DGNN layers.
+  • Changes: Compared with the baseline AT-DGNN-Gen.
+As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of critical components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and the dynamic graph neural network (D).
+The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
+Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance.
+Modifying the number of DGNN layers to $1$, $2$, or $4$ led to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively.
+A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
+The optimal configuration, employing $3$ layers of the DGNN, is found to be the most effective, as evidenced by achieving the highest scores in both ACC and F1 metrics.
+These findings substantiate the design choices of the AT-DGNN model and shed light on the interaction among its components, significantly enhancing computational neuroscience analysis.
 VI Conclusion
-This study introduces the MEEG dataset and the AT-DGNN framework as significant advancements in EEG-based emotion recognition.
-The MEEG dataset, which uses music to induce emotional states, provides a valuable resource for studying brain responses to emotional stimuli.
-The AT-DGNN framework, integrating attention mechanisms with DGNN, excels in capturing complex temporal brain dynamics, significantly improving emotion recognition accuracy compared to other SOTA methods.
-These innovations advance brain-computer interface technologies and open new avenues for exploring the relationship between music, emotion, and brain activity.
-The superior performance of AT-DGNN, validated against traditional models, underscores its potential for enhancing emotion analysis.
+This study introduces the MEEG dataset and the AT-DGNN framework, marking substantial progress in the field of EEG-based emotion recognition.
+The MEEG dataset uniquely employs music to induce emotional states, providing an invaluable resource for probing brain responses to emotional stimuli.
+The AT-DGNN, which integrates attention mechanisms with DGNN, adeptly captures the intricate temporal dynamics of brain activity, thereby markedly enhancing emotion recognition accuracy over existing SOTA methods.
+This advancement not only propels BCI technology forward but also facilitates novel investigations into the interplay among music, emotion, and cerebral activity.
+The exemplary performance of AT-DGNN, validated through rigorous comparisons with traditional models, solidifies its utility in advancing neuropsychological understanding.

### v3

@@
 Abstract
-Neuropsychological research highlights the essential role of coordinated activities across brain regions during cognitive tasks.
-This study expands the existing EEG datasets by constructing the MEEG dataset, a multi-modal compilation of music-induced electroencephalogram (EEG) recordings, and further investigates the brain network topology during emotional responses to music.
-The MEEG dataset, capturing emotional responses to various musical stimuli across different valence and arousal levels, enables an in-depth analysis of brainwave patterns within musical contexts.
-We introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG emotion recognition.
-By integrating an attention mechanism with a dynamic graph neural network (DGNN), the AT-DGNN model captures complex local and global EEG dynamics, demonstrating superior performance with accuracy of 83.74% in arousal and 86.01% in valence, outperforming current state-of-the-art (SOTA) methods.
-Comparative analyses with traditional datasets like DEAP underscore our model’s effectiveness, highlighting the potential of music as a potent emotional stimulus.
-This study advances graph-based learning techniques in brain-computer interfaces (BCI), significantly enhancing the precision of EEG-based emotion recognition and deepening our understanding of cognitive functions in various brain regions.
-The source code and dataset are accessible at https://github.com/xmh1011/AT-DGNN.
+We present the MEEG dataset, a multi-modal collection of music-induced electroencephalogram (EEG) recordings designed to capture emotional responses to various musical stimuli across different valence and arousal levels.
+This public dataset facilitates an in-depth examination of brainwave patterns within musical contexts, providing a robust foundation for studying brain network topology during emotional processing.
+Leveraging the MEEG dataset, we introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition.
+This model combines an attention mechanism with a dynamic graph neural network (DGNN) to capture intricate EEG dynamics.
+The AT-DGNN achieves state-of-the-art (SOTA) performance with an accuracy of 83.74% in arousal recognition and 86.01% in valence recognition, outperforming existing SOTA methods.
+Comparative analysis with traditional datasets, such as DEAP, further validates the model’s effectiveness and underscores the potency of music as an emotional stimulus.
+This study advances graph-based learning methodology in brain-computer interfaces (BCI), significantly improving the accuracy of EEG-based emotion recognition.
+The MEEG dataset and source code are publicly available at https://github.com/xmh1011/AT-DGNN.
 Index Terms—electroencephalogram, emotion recognition, dynamic graph neural networks.
@@
 The model conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
-BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-machine interaction [3, 4].
+BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-computer interaction [3].
 This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands.
@@
 In 2023, Ding et al.
-introduced the Local-Global-Graph network (LGGNet) to process EEG data in an image-like format, improving classification performance [16].
-However, LGGNet’s single-layer graph neural network (GNN) struggles to capture the dynamic features of EEG signals.
-Additionally, popular EEG datasets like DEAP and SEED primarily use video stimuli, limiting the study of the relationship between music, emotions, and EEG signals.
-In contrast, EEG datasets and research using other types of stimuli are relatively scarce.
-To address shortcomings in emotion recognition research, we create the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format with music-induced emotional states to enhance ecological validity.
-We improve the LGGNet architecture by integrating an attention mechanism and using a multi-layer dynamic graph neural network, significantly boosting classification accuracy on the MEEG dataset.
-This innovative approach allows for a more nuanced examination of brain function interconnections via graph-based learning.
+introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, which improved classification performance [15].
+Alternatively, LGGNet’s single-layer graph neural network faces challenges in capturing the dynamic features of EEG signals.
+Additionally, widely used EEG datasets such as DEAP and SEED primarily involve visual stimuli, which limits exploration of the relationship between music, emotions, and EEG signals.
+Conversely, EEG datasets and research using musical stimuli are relatively scarce.
+Moreover, existing SOTA methods achieve lower accuracy on the DEAP dataset, thus limiting comprehensive performance evaluation of models.
+To address these limitations in emotion recognition research, the MEEG (Music EEG) dataset is developed accordingly.
+This multi-modal dataset, similar to DEAP, utilizes music-induced emotional states to improve the accuracy of emotion recognition.
+By integrating a sliding window technique, an attention mechanism, and a multi-layer dynamic graph neural network into the LGGNet architecture, significant improvements in classification accuracy are achieved.
 The contributions of our study are threefold and can be summarized as follows:
   1. 1.
-    We establish the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, enhanced with various types of music to effectively induce emotional states.
-Experimental results show that the MEEG dataset outperforms the video-stimuli-based DEAP dataset in emotion induction, effectively enhancing model ACC and F1 scores.
+    The MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, is enhanced with diverse music to induce emotional states effectively.
+It outperforms the DEAP dataset in emotion induction, improving model accuracy (ACC) and F1 scores.
   2. 2.
-    We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain.
-By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion recognition.
+    The AT-DGNN framework explores connections within and between brain functional areas.
+By integrating a sliding window technique, an attention mechanism, and a stacked DGNN, this novel architecture for emotion recognition enhances the learning of dynamic features in EEG recordings.
   3. 3.
-    We compare the performance of the proposed method with CNN, TCN, and GNN-based SOTA methods on the MEEG dataset. Extensive ablation experiments are conducted to better understand AT-DGNN.
+    Performance of the AT-DGNN is compared with other CNN, TCN, and GNN-based SOTA methods on MEEG dataset. Ablation experiments provide insights into the AT-DGNN framework.
 II Related Work
@@
 The initial sampling rate of the MEEG dataset is 1000 Hz.
-To align with standard EEG datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz.
+To align with standard datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz.
 This preprocessing phase involves the use of a band-pass filter with a range of 1 to 50 Hz to enhance the signal quality by minimizing interference.
-Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23].
+Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz).
 III-B Feature Extraction
@@
 These kernels obviate the need for manually extracted features.
-The $i$th-kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), is given by [24]:
+The $i$-th kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), can be represented by [22]:
 $S_{T}^{i}=(1,\alpha^{i}\cdot f_{s}),\quad i\in[1,2,3]$ (5)
@@
 III-B2 Sliding Window Segmentation
-Following the temporal learner, a convolution-based sliding window technique as [25] is applied to segment the EEG time series $X$ into multiple windows.
+Following the temporal learner, a convolution-based sliding window technique as described in [23] is applied to segment the EEG time series $X$ into multiple windows.
 This method integrates sliding window segmentation with convolutional operations, significantly reducing computational overhead by allowing convolutions to be executed once across all windows.
-This approach not only enhances data augmentation but also accelerates processing through parallel computation.
+This approach enhances data augmentation while accelerating processing through parallel computation.
 The time series $X$ is segmented into windows $X_{w}\in\mathbb{R}^{B\times C\times W}$ using a window of length $W$ and stride $S$, where $w=1,\ldots,n$ represents the window index and $n$ is the total number of windows.
@@
 The normalized EEG data are denoted as $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ represents the number of EEG channels and $l$ represents the temporal dimension’s sample length.
-To extract dynamic temporal features, we employ a temporal convolutional block on each window of the EEG data.
+To extract dynamic temporal features, a temporal convolutional block is applied to each window of the EEG data.
 The normalized EEG data $X_{i}$ serves as the input to this block.
-The temporal convolutional block processes $X_{i}$ and generates an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ represents the window’s temporal length and $w$ indexes the window.
-Batch normalization is applied to the output to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity.
+The temporal convolutional block processes $X_{i}$ to produce an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ is the temporal length of the window and $w$ is the window index.
+Batch normalization is applied to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity.
 The output $Z_{\textit{tcn}}^{w}$ is computed as follows:
@@
 III-C1 Graph Filtering Layer
-For the extracted features, we adopt the method described by [16], which involves defining functional areas and performing local filtering.
-The electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig.
+The method described by [15] is adopted for the extracted features, involving the definition of functional areas and local filtering.
+Electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig.
 2.
-Each channel is treated as a node, and the learned dynamic temporal representations are considered as node attributes.
-Figure 2: Local global graph definitions [16]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
+Each channel is treated as a node, with the learned dynamic temporal representations considered as node attributes.
+Figure 2: Local-global graph definitions [15]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
 In the preprocessing step, EEG channels are systematically reordered within predefined groups to ensure adjacency of channels within each local graph.
@@
 20$\triangleright$ Graph-based Learning
-21:do graph filtering and aggregation on each node by (12) - 15 to get $Z^{i}_{\textit{local}}$;
+21:perform graph filtering and aggregation on each node by (12) - (15) to get $Z^{i}_{\textit{local}}$;
 22:for $i\leftarrow 1$ to $3$ do
@@
 26:Return pred
-We introduce a stacked DGNN architecture to discern complex structural relationships within graphs.
-Each layer adaptively recalculates the adjacency matrix based on input features, effectively capturing hierarchical feature variations.
+A stacked Directed Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
 The architecture incorporates multiple GNN layers.
@@
 Considering that the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [25], we opted to induce emotions in participants using music in the MEEG dataset.
-We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral.
+Under the guidance of music professors, we selected 20 lesser-known Western classical music clips by composers such as Shostakovich and Tchaikovsky.
+Each clip lasts one minute and is characterized by distinct levels of valence and arousal.
 To minimize emotional carryover, a 15-second pause was introduced between clips [26].
-The chosen music pieces included lesser-known works of Western classical music by composers such as Shostakovich and Tchaikovsky.
 The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music.
@@
 The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.
+In contrast to the DEAP dataset, the MEEG dataset reduces experimental bias by rigorously selecting participants and musical stimuli, minimizing subjective factors and improving the effectiveness of emotional elicitation.
 IV-B Experiment Settings
@@
 Initially, the best model identified from the $k$-fold cross-validation is saved as a preliminary candidate.
-This model is then refined using combined training data from the $k$ folds, fine-tuned at a reduced learning rate to prevent overfitting, and further trained for up to 20 epochs or until it achieves a training accuracy of 100%, ensuring precise calibration.
-Importantly, test data is excluded from the training process to maintain the integrity of the evaluation.
+This model is subsequently refined with the aggregated data from all $k$ folds, fine-tuned at a lower learning rate to avoid overfitting, and further trained for a maximum of 20 epochs or until it reaches 100% training accuracy, ensuring precise calibration.
+Importantly, test data is excluded from the training to preserve evaluation integrity.
 The integration of two-stage training with $10$-fold cross-validation minimizes variability in assessment results, providing a comprehensive and reliable evaluation framework suitable for diverse research and application domains.
 IV-C Implement Details
-The model was implemented using PyTorch [30] library, and the code can be found at https://github.com/xmh1011/AT-DGNN.
+The model was implemented using PyTorch [28] library.
 Cross-entropy loss was chosen as the objective function to guide the training process.
 The training was divided into two stages, with the first stage capped at 200 epochs and the second at 20 epochs.
-The batch size was set to 64.
 To reduce training time and prevent overfitting, early stopping was implemented.
-For the attention module, the window size was set to half the data frequency.
+For the attention module, the window size was set to half the $f_{s}$.
 The kernel sizes for the time learner were set to 100, 50, and 25.
-The hidden layer count for the DGNN was established at 32.
 Training was optimized using the Adam optimizer, starting with an initial learning rate of $31e-$ 3, which was reduced by a factor of 10 during the second stage.
-For more details, please refer to the open-source GitHub repository for AT-DGNN.
+For more information on the data processing, model implementation details, and accessing the dataset, please refer to our GitHub repository https://github.com/xmh1011/AT-DGNN.
 TABLE I: Comparison of ACC and F1 scores (mean $\pm$ std) for the MEEG dataset under trial-wise $10$-fold cross-validation
@@
 DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44 82.42 82.12
-LGGNet-Fro 81.85 12.57 81.58 12.73 82.34 12.84 81.66 13.91 82.11 81.62
-LGGNet-Gen 82.15 12.72 82.15 12.57 82.53 13.20 82.47 13.22 82.34 82.31
-LGGNet-Hem 81.92 12.08 81.58 12.43 82.37 12.71 81.58 14.13 82.15 81.58
+LGGNet-Fro 81.85 12.57 81.58 12.73 84.28 11.83 83.87 12.66 83.07 82.73
+LGGNet-Gen 82.15 12.72 82.15 12.57 84.53 11.44 84.38 11.56 83.34 83.27
+LGGNet-Hem 81.92 12.08 81.58 12.43 84.93 11.21 84.46 12.27 83.42 83.02
 AT-DGNN-Fro 83.51 11.09 83.06 11.83 85.56 10.69 85.61 10.89 84.54 84.34
@@
 As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy.
-The AT-DGNN-Gen model not only achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, but also records the highest accuracy of 86.01% in the valence dimension, a substantial improvement of 3.48% over the LGGNet series.
+The AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and also records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series.
 This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
@@
 This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks.
-Furthermore, consistent with the findings of [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
-Furthermore, these results demonstrate a marked improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy surpassing previous models by up to 7.16% and 5.78%, respectively.
-This emphasizes the advanced capability of the DGNN approach in graph-structured data analysis, providing more accurate and reliable emotion recognition.
+Furthermore, consistent with the findings of LGGNet [15], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
+Furthermore, the results show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively.
+This highlights the enhanced capability of the DGNN approach in analyzing graph-structured data, offering more accurate and reliable emotion recognition.
+Additionally, the AT-DGNN exhibits a lower standard deviation compared to other models, indicating more balanced performance across various samples and superior generalization ability.
 All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%.
-On the DEAP dataset, AT-DGNN averages 60.56% accuracy across four dimensions.
+On the DEAP dataset, AT-DGNN achieves an average accuracy of 60.56% across four dimensions, while LGGNet reaches 60.96%.
 According to Huang et al.
-[32], other models, including LGGNet, also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not.
-All models perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
+[29], other SOTA methods also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not.
+All methods perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
 V-B Ablation Study
@@
   • Changes: Compared with the baseline AT-DGNN-Gen.
-As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of critical components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and the dynamic graph neural network (D).
+As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of key components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and dynamic graph neural network (D).
 The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
 Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance.
-Modifying the number of DGNN layers to $1$, $2$, or $4$ led to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively.
+Modifying the number of DGNN layers to $1$, $2$, or $4$ lead to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively.
 A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
-The optimal configuration, employing $3$ layers of the DGNN, is found to be the most effective, as evidenced by achieving the highest scores in both ACC and F1 metrics.
-These findings substantiate the design choices of the AT-DGNN model and shed light on the interaction among its components, significantly enhancing computational neuroscience analysis.
+The optimal configuration, utilizing $3$ layers of the DGNN, demonstrated superior effectiveness by achieving the highest scores in both ACC and F1 metrics, substantiating the design choices and highlighting the critical interdependence of its three main components: sliding window segmentation, multi-head attention mechanisms, and dynamic graph neural networks.
+These findings reveal how the synergistic operation of these elements is essential for capturing the dynamic temporal features of EEG data, significantly enhancing the model’s capability to process EEG data effectively and advancing computational neuroscience analysis.
 VI Conclusion
-This study introduces the MEEG dataset and the AT-DGNN framework, marking substantial progress in the field of EEG-based emotion recognition.
-The MEEG dataset uniquely employs music to induce emotional states, providing an invaluable resource for probing brain responses to emotional stimuli.
-The AT-DGNN, which integrates attention mechanisms with DGNN, adeptly captures the intricate temporal dynamics of brain activity, thereby markedly enhancing emotion recognition accuracy over existing SOTA methods.
-This advancement not only propels BCI technology forward but also facilitates novel investigations into the interplay among music, emotion, and cerebral activity.
-The exemplary performance of AT-DGNN, validated through rigorous comparisons with traditional models, solidifies its utility in advancing neuropsychological understanding.
+This study introduces the MEEG dataset and the AT-DGNN framework, significantly advancing EEG-based emotion recognition.
+The MEEG dataset utilizes music to induce emotional states, providing a unique and critical resource for analyzing brain responses to emotional stimuli.
+The AT-DGNN framework captures the complex temporal dynamics of brain activity, enhancing emotion recognition accuracy beyond existing SOTA methods.
+Moreover, the MEEG dataset displays emotional states of subjects with greater precision, enabling SOTA methods to achieve higher accuracy and establishing a more robust benchmark for evaluating EEG-based emotion analysis models.
+These advancements propel BCI technology forward and facilitate new research into the relationships between music, emotion, and brain activity.
+VII ACKNOWLEDGEMENTS
+This paper is supported in part by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009.

### v4

@@
 The AT-DGNN achieves state-of-the-art (SOTA) performance with an accuracy of 83.74% in arousal recognition and 86.01% in valence recognition, outperforming existing SOTA methods.
-Comparative analysis with traditional datasets, such as DEAP, further validates the model’s effectiveness and underscores the potency of music as an emotional stimulus.
 This study advances graph-based learning methodology in brain-computer interfaces (BCI), significantly improving the accuracy of EEG-based emotion recognition.
@@
 I Introduction
-Emotion, a complex physiological and psychological phenomenon, plays a critical role in interpreting genuine reactions to daily interactions [1].
-In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2].
-The model conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
-BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-computer interaction [3].
-This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands.
-Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [4, 5].
+Emotion plays a key role in interpreting human responses, and Russell’s Valence-Arousal (VA) model offers a dimensional framework for quantifying emotional states within a continuous space [1].
+Meanwhile, BCI technology, which decodes EEG signals reflecting cortical voltage fluctuations, translates thoughts into commands for external devices, significantly enhancing human-computer interaction [2].
+With their high temporal resolution, objectivity, affordability, and quick acquisition, EEG signals are crucial for emotion recognition and neuroscience research, particularly in integrating emotional models with advanced decoding methods [3].
 In 2011, Koelstra et al.
-established the DEAP dataset, which consists of physiological signals, such as EEG, elicited by emotional responses to musical videos [6].
-Subsequently, Zheng et al.
-introduced the SEED dataset in 2015, comprising well-annotated EEG signals induced by cinematic stimuli, with data collected from multiple subjects [7].
-These datasets have significantly contributed to the advancement of computational models for emotion analysis.
-Over the past decades, a variety of machine learning and signal processing techniques have been applied to emotion recognition based on physiological data.
+introduced the DEAP dataset, consisting of physiological signals like EEG elicited by emotional responses to musical videos [4], followed by Zheng et al.’s SEED dataset in 2015, which provided well-annotated EEG signals from cinematic stimuli across multiple subjects [5].
+These datasets have greatly advanced computational models for emotion analysis.
 Recent advances in deep learning have significantly impacted EEG signal processing.
@@
 This assertion was corroborated by Musallam et al., who demonstrated the versatility and efficiency of TCNs in complex BCI tasks, particularly through the integration of TCNs in motor imagery classification [9].
-All the aforementioned research treated EEG signals primarily as two-dimensional time-series data, with dimensions corresponding to channels and time [12].
-Here, channels denote EEG electrodes, positioned on the scalp using the 10-20 system, as prescribed by the International Federation of Clinical Neurophysiology, to capture neural activity from distinct brain regions [3].
-More recently, a paradigm shift has occurred with EEG data now being conceptualized as graphs, representing the electrodes’ two-dimensional arrangement on the scalp [13].
-In this model, raw data or extracted features from each electrode augment this spatial framework by adding a third dimension [14].
-In 2023, Ding et al.
-introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, which improved classification performance [15].
+All aforementioned studies treated EEG signals as two-dimensional time-series data, with channels representing EEG electrodes positioned according to the 10-20 system to capture neural activity across brain regions [10, 2].
+Recently, EEG data have been increasingly modeled as graphs, representing the spatial arrangement of electrodes, where features from each electrode add a third dimension to the model [11, 12].
+In 2019, Song et al.
+introduced a novel DGCNN for multichannel EEG emotion recognition, utilizing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination, demonstrating superiority over existing approaches [13].
+Subsequently, advancements in GNNs have been increasingly applied to emotion recognition.
+Bao et al.
+integrated a multi-layer GNN with a style-reconfigurable CNN [14], while Asadzadeh et al.
+further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance [15].
+Building on prior advancements, greater progress was made in 2023 when Ding et al.
+introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, achieving improved classification performance [16].
 Alternatively, LGGNet’s single-layer graph neural network faces challenges in capturing the dynamic features of EEG signals.
@@
 II-A Multi-head Attention Mechanism
-EEG data is inherently complex, characterized by temporal dependencies and discrete features.
-Utilizing multiple attention heads allows us to capture a diverse range of temporal patterns and dependencies, effectively identifying key temporal features.
-In this module, the attention block consists of a multi-head attention (MHA) layer with several self-attention heads.
+EEG data is inherently complex, with significant temporal dependencies and discrete features.
+Multi-head attention (MHA) helps capture diverse temporal patterns and dependencies, enabling the identification of key features.
+In this module, the attention block consists of a multi-head attention layer with several self-attention heads.
+As shown by Zhang et al.
+[17], multi-head self-attention effectively integrates multi-source information, improving the model’s robustness in handling complex temporal features in EEG data, leading to enhanced classification performance.
 Each self-attention head comprises three fundamental components: queries $Q$, keys $K$, and values $V$.
@@
 Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma(\cdot)$ being a non-linear activation function.
-II-C Graph Neural Networks for EEG
-In 2019, Song et al.
-[17] first introduced a novel DGCNN specifically for multichannel EEG emotion recognition, employing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination.
-This method proved to be superior to existing approaches.
-Subsequently, advancements in GNNs have been increasingly applied to emotion recognition.
-On the SEED dataset, Bao et al.
-[18] integrated a multi-layer GNN with a style-reconfigurable CNN, while Asadzadeh et al.
-[19] further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance.
-For the DEAP dataset, Zhao et al.
-[20] developed coherence-based node clustering and multi-pooling techniques for GCNs.
-Additionally, Zhang et al.
-[21] enhanced GCN feature extraction by using Granger causality analysis and differential entropy to construct more informative connectivity matrices.
 III Methodology
@@
 III-C2 Stacked DGNN
+A stacked Dynamic Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
+The architecture incorporates multiple GNN layers.
+In each layer, features are grouped and aggregated into matrices, resulting in an aggregated feature matrix $Z_{\textit{agg}}$.
+The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$:
+$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
+Self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is normalized as:
+$A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (17)
+Here, $\tilde{D}$ is the diagonal degree matrix of $A$, with each diagonal element $\tilde{D}_{ii}$ being the sum of the $i$-th row of $A$.
+During forward propagation, each DGNN layer processes the inputs using a normalized adjacency matrix $\tilde{A}_{i}$, a weight matrix $W_{i}$, and a bias vector $b_{i}$ for graph convolution.
+The outputs for each layer are calculated as:
+$\begin{cases}H_{1}=\textit{ReLU}(\tilde{A}_{1}\cdot Z_{\textit{agg}}\cdot W_{1}+b_{1})\\ H_{i}=\textit{ReLU}(\tilde{A}_{i}\cdot H_{i-1}\cdot W_{i}+b_{i}),\quad i\in[2,n-1]\\ H_{n}=\textit{ReLU}(\tilde{A}_{n}\cdot H_{n-1}\cdot W_{n}+b_{n})\end{cases}$ (18)
+Here, $n$ denotes the total number of layers in the DGNN, which is set to $3$ in our study.
+The final Output is computed through a sequence of operations including batch normalization $\mathcal{F}_{\textit{bn}}(\cdot)$, dropout $\mathcal{F}_{\textit{dropout}}(\cdot)$, and softmax activation $\Phi_{\textit{softmax}}(\cdot)$, structured as:
+$\textit{Output}=\Phi_{\textit{softmax}}(\mathcal{F}_{\textit{dropout}}(\Gamma(\mathcal{F}_{\textit{bn}}(H^{(L)}))))$ (19)
+where $\Gamma(\cdot)$ denotes the flattening operation, ensuring that the network output is appropriately structured for subsequent processing or analysis.
+Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
 Algorithm 1 AT-DGNN
@@
 26:Return pred
-A stacked Directed Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
-The architecture incorporates multiple GNN layers.
-In each layer, features are grouped and aggregated into matrices, resulting in an aggregated feature matrix $Z_{\textit{agg}}$.
-The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$:
-$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
-Self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is normalized as:
-$A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (17)
-Here, $\tilde{D}$ is the diagonal degree matrix of $A$, with each diagonal element $\tilde{D}_{ii}$ being the sum of the $i$-th row of $A$.
-During forward propagation, each DGNN layer processes the inputs using a normalized adjacency matrix $\tilde{A}_{i}$, a weight matrix $W_{i}$, and a bias vector $b_{i}$ for graph convolution.
-The outputs for each layer are calculated as:
-$\begin{cases}H_{1}=\textit{ReLU}(\tilde{A}_{1}\cdot Z_{\textit{agg}}\cdot W_{1}+b_{1})\\ H_{i}=\textit{ReLU}(\tilde{A}_{i}\cdot H_{i-1}\cdot W_{i}+b_{i}),\quad i\in[2,n-1]\\ H_{n}=\textit{ReLU}(\tilde{A}_{n}\cdot H_{n-1}\cdot W_{n}+b_{n})\end{cases}$ (18)
-Here, $n$ denotes the total number of layers in the DGNN, which is set to $3$ in our study.
-The final Output is computed through a sequence of operations including batch normalization $\mathcal{F}_{\textit{bn}}(\cdot)$, dropout $\mathcal{F}_{\textit{dropout}}(\cdot)$, and softmax activation $\Phi_{\textit{softmax}}(\cdot)$, structured as:
-$\textit{Output}=\Phi_{\textit{softmax}}(\mathcal{F}_{\textit{dropout}}(\Gamma(\mathcal{F}_{\textit{bn}}(H^{(L)}))))$ (19)
-where $\Gamma(\cdot)$ denotes the flattening operation, ensuring that the network output is appropriately structured for subsequent processing or analysis.
-Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
 IV Experiments
@@
 For more information on the data processing, model implementation details, and accessing the dataset, please refer to our GitHub repository https://github.com/xmh1011/AT-DGNN.
-TABLE I: Comparison of ACC and F1 scores (mean $\pm$ std) for the MEEG dataset under trial-wise $10$-fold cross-validation
-Method Arousal Valence Emotion
-ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%) ACC (%) F1 (%)
-EEGNet 75.47 17.63 75.54 17.83 73.66 16.82 73.33 16.94 74.57 74.44
-DeepConvNet 79.35 15.03 78.43 18.10 72.38 18.87 73.03 19.27 75.87 75.73
-ShallowConvNet 80.96 14.49 80.87 14.71 79.77 16.44 80.22 15.60 80.37 80.55
-TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08 78.79 81.24
-EEG-TCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43 73.04 64.04
-TCNet-Fusion 76.72 17.32 76.95 17.49 75.88 16.70 75.42 16.97 76.30 76,20
-ATCNet 82.01 14.06 79.17 17.57 83.19 12.09 82.08 13.41 82.60 80.63
-DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44 82.42 82.12
-LGGNet-Fro 81.85 12.57 81.58 12.73 84.28 11.83 83.87 12.66 83.07 82.73
-LGGNet-Gen 82.15 12.72 82.15 12.57 84.53 11.44 84.38 11.56 83.34 83.27
-LGGNet-Hem 81.92 12.08 81.58 12.43 84.93 11.21 84.46 12.27 83.42 83.02
-AT-DGNN-Fro 83.51 11.09 83.06 11.83 85.56 10.69 85.61 10.89 84.54 84.34
-AT-DGNN-Gen 83.74 12.03 84.56 10.88 86.01 9.40 85.68 10.06 84.88 85.12
-AT-DGNN-Hem 83.73 11.20 84.78 10.93 84.89 10.57 85.35 10.24 84.31 85.07
+TABLE I: Comparison of ACC and F1 scores for the MEEG dataset and DEAP dataset under trial-wise 10-fold cross-validation
+Method MEEG Arousal MEEG Valence MEEG Emotion DEAP Emotion
+ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%)
+EEGNet 75.47∗∗ 75.54∗∗ 73.66∗∗ 73.33∗∗ 74.57 74.44 60.36 64.64
+DeepConvNet 79.35∗∗ 78.43∗∗ 72.38∗∗ 73.03∗∗ 75.87 75.73 60.41 63.80
+ShallowConvNet 80.96∗∗ 80.87∗∗ 79.77∗∗ 80.22∗∗ 80.37 80.55 61.01 64.44
+TSception 79.23∗∗ 81.30 78.34∗∗ 81.18∗∗ 78.79 81.24 60.42 64.15
+EEG-TCNet 76.36∗∗ 73.79∗∗ 69.72∗∗ 54.28∗∗∗ 73.04 64.04 56.11 52.50
+TCNet-Fusion 76.72∗∗ 76.95∗∗ 75.88∗∗ 75.42∗∗ 76.30 76.20 59.33 63.94
+ATCNet 82.01∗∗ 79.17∗∗ 83.19∗∗ 82.08∗∗ 82.60 80.63 60.83 62.32
+DGCNN 82.12∗∗ 81.83∗∗ 82.72∗∗ 82.40∗∗ 82.42 82.12 60.36 61.27
+LGGNet-Fro 81.85∗∗ 81.58∗∗ 84.28∗∗ 83.87∗∗ 83.07 82.73 60.59 64.48
+LGGNet-Gen 82.15∗∗ 82.15∗∗ 84.53∗∗ 84.38∗∗ 83.34 83.27 60.96 65.41
+LGGNet-Hem 81.92∗∗ 81.58∗∗ 84.93∗∗ 84.46∗∗ 83.42 83.02 59.88 62.50
+AT-DGNN-Fro 83.51∗∗ 83.06∗∗ 85.56∗∗ 85.61∗∗ 84.54 84.34 59.95 62.61
+AT-DGNN-Gen 83.74∗∗ 84.56∗∗ 86.01 85.68 84.88 85.12 60.55 63.83
+AT-DGNN-Hem 83.73∗∗ 84.78∗∗ 84.89∗∗ 85.35∗∗ 84.31 85.07 59.36 62.16
+  • * indicating $std<10\%$, ** indicating $std<20\%$, *** indicating $std<30\%$.
   • Emotion averages the ACC and F1 scores of both the arousal and valence dimensions.
-  • The results for all models are derived from the optimal values among ten trials with distinct random seeds.
+  • The results for are derived from the optimal values among ten trials with distinct random seeds.
 V RESULTS AND DISCUSSION
@@
 V-A Emotion Recognition
-As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy.
-The AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and also records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series.
+As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous state-of-the-art LGGNet series in both arousal and valence dimensions, with notable increases in F1 scores and accuracy.
+Specifically, the AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series.
 This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
-The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions.
-This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks.
+Moreover, the AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions.
+This highlights the effectiveness of our graph-based approach in handling complex emotion recognition tasks.
 Furthermore, consistent with the findings of LGGNet [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
-Furthermore, the results show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively.
+The results also show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively.
 This highlights the enhanced capability of the DGNN approach in analyzing graph-structured data, offering more accurate and reliable emotion recognition.
 Additionally, the AT-DGNN exhibits a lower standard deviation compared to other models, indicating more balanced performance across various samples and superior generalization ability.
-All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%.
-On the DEAP dataset, AT-DGNN achieves an average accuracy of 60.56% across four dimensions, while LGGNet reaches 60.96%.
-According to Huang et al.
-[29], other SOTA methods also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not.
-All methods perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
+On both the MEEG and DEAP datasets, the AT-DGNN demonstrates strong generalization ability.
+All models achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%.
+Specifically, all models, including AT-DGNN, achieve significantly higher accuracy on the MEEG dataset than on the DEAP dataset, suggesting that music is more effective than video in evoking emotions and indicating the higher quality of the MEEG dataset.
+On the DEAP dataset, the AT-DGNN achieves an average accuracy of 60.56% across four dimensions, close to other SOTA models and surpassing some of them, which demonstrates the AT-DGNN model’s strong generalization capability across different datasets.
 V-B Ablation Study
 TABLE II: Ablation study on Emotion of MEEG using AT-DGNN-Gen.
-S M D ACC(%) Changes(%) F1(%) Changes(%)
-81.91 -2.97 81.81 -3.31
-3 78.31 -6.57 79.46 -5.66
-3 80.89 -3.99 81.16 -3.96
-1 81.91 -2.97 81.81 -3.31
-2 82.94 -1.94 82.52 -2.60
-4 83.17 -1.71 83.48 -1.64
-3 84.88 - 85.12 -
+S A G T ACC(%) Changes(%) F1(%) Changes(%)
+3 3 78.31 -6.57 79.46 -5.66
+3 3 80.89 -3.99 81.16 -3.96
+1 3 81.91 -2.97 81.81 -3.31
+2 3 82.94 -1.94 82.52 -2.60
+4 3 83.17 -1.71 83.48 -1.64
+3 83.05 -1.83 83.89 -1.23
+3 1 82.95 -1.93 82.84 -2.28
+3 2 82.60 -2.28 81.91 -3.21
+3 4 82.89 -1.99 83.43 -1.69
+3 74.64 -10.24 74.84 -10.28
+3 3 84.88 - 85.12 -
   • S: Sliding window segmentation.
-  • M: Multi-head attention mechanism.
-  • D: Dynamic graph neural network.
+  • A: Multi-head attention mechanism.
+  • G: Graph neural network.
+  • T: Temporal learner.
   • ✓: Keep the component.
   • ✗: Exclude the component.
-  • 1, 2, 3, 4: The number of stacked DGNN layers.
+  • 1, 2, 3, 4: The number of DGNN layers or temporal learner.
   • Changes: Compared with the baseline AT-DGNN-Gen.
-As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of key components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and dynamic graph neural network (D).
-The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
-Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance.
-Modifying the number of DGNN layers to $1$, $2$, or $4$ lead to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively.
-A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
-The optimal configuration, utilizing $3$ layers of the DGNN, demonstrated superior effectiveness by achieving the highest scores in both ACC and F1 metrics, substantiating the design choices and highlighting the critical interdependence of its three main components: sliding window segmentation, multi-head attention mechanisms, and dynamic graph neural networks.
-These findings reveal how the synergistic operation of these elements is essential for capturing the dynamic temporal features of EEG data, significantly enhancing the model’s capability to process EEG data effectively and advancing computational neuroscience analysis.
+As shown in Table II, the ablation study systematically evaluates the contributions of key components in the AT-DGNN-Gen model, including sliding window segmentation (S), multi-head attention mechanism (A), graph neural network layers (G), and temporal learner layers (T).
+The baseline model, incorporating all components with three layers each for G and T, achieved the highest ACC of 84.88% and F1 score of 85.12%.
+Excluding the multi-head attention mechanism (A) caused a significant performance drop, with ACC decreasing by 6.57% and F1 score by 5.66%.
+Similarly, removing the sliding window segmentation (S) resulted in reductions of 3.99% in ACC and 3.96% in F1 score.
+These results confirm the essential roles of S and A in feature extraction and temporal modeling.
+Varying the number of G layers showed that using three layers yielded optimal results, while deviations caused performance declines.
+Increasing G layers to four resulted in a minor ACC drop of 1.71%, while reducing to one layer led to a larger decline of 2.97%.
+Similar trends were observed for T layers, where three layers provided the best balance, and exclusion of T caused a substantial drop of 10.24% in ACC and 10.28% in F1 score, highlighting its critical role in capturing temporal dynamics.
+Overall, the ablation study demonstrates that each component of the AT-DGNN-Gen model contributes significantly to its performance.
+The sliding window segmentation and multi-head attention mechanism are crucial for effective feature extraction, while the optimal number of layers for the graph neural network and temporal learner is three, balancing model complexity and performance.
+These findings substantiate the design choices of our model and underscore the importance of the synergistic operation of its components in enhancing emotion recognition from EEG data.
+The results emphasize how the model’s architecture is essential for capturing the complex spatiotemporal patterns inherent in EEG signals, thereby advancing computational neuroscience analysis.
 VI Conclusion
@@
 VII ACKNOWLEDGEMENTS
-This paper is supported in part by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009.
+This paper is supported by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009. and Key Lab of Information Network Security, Ministry of Public Security.